### v1

# DirectMultiStep: Direct Route Generation for Multi-Step Retrosynthesis

Abstract
Traditional computer-aided synthesis planning (CASP) methods rely on iterative single-step predictions, leading to exponential search space growth that limits efficiency and scalability. We introduce a transformer-based model that directly generates multi-step synthetic routes as a single string by conditionally predicting each molecule based on all preceding ones. The model accommodates specific conditions such as the desired number of steps and starting materials, outperforming state-of-the-art methods on the PaRoutes dataset with a 2.2x improvement in Top-1 accuracy on the n1 test set and a 3.3x improvement on the n5 test set. It also successfully predicts routes for FDA-approved drugs not included in the training data, showcasing its generalization capabilities. While the current suboptimal diversity of the training set may impact performance on less common reaction types, our approach presents a promising direction towards fully automated retrosynthetic planning.
1 Introduction
Finding the most efficient route to a desired chemical compound is a daily challenge for synthetic organic chemists in both fundamental research and drug discovery. Route efficiency is determined by various factors, some of which can be objectively assessed, such as overall yield (not all chemical reactions have 100% conversion rate) and enantiomeric excess (in case of chiral compounds), where higher values are always preferred. Other factors, such as atom efficiency (minimization of byproducts) and availability (cost) of starting materials, are more case-dependent. A used chemical reactant is considered waste unless it can be repurposed as a reactant in a different process. Similarly, the choice of starting materials depends on factors such as budget, logistics, and the availability of specific equipment. It’s worth noting that many commercially available compounds can be synthesized from other commercially available compounds, adding another layer of complexity to the decision-making process.
A significant contribution to the development of algorithmic frameworks for identifying synthetic routes leading to any desired compound (subject of the 1990 Nobel Prize in Chemistry) was made by Elias James Corey. Corey’s framework, now known as retrosynthetic analysis, begins with identifying atoms that would serve as reaction centers. Disconnecting bonds between these centers results in the formation of hypothetical fragments (called synthons) from which a precursor molecule can be created. This mapping from fragments to actual molecules is one-to-many because there is usually more than one functional group that could participate in any given type of reaction. Importantly, meticulous application of Corey’s framework (i.e., systematically breaking small subsets of bonds) will eventually lead to commercially available starting materials. The algorithmic nature of this process allowed Corey to envision automating these rules to create Computer-Aided Synthesis Planning (CASP) as early as 1969 [1].
Recent advancements in data science and machine learning (ML) methods have led to a surge of interest in developing CASP methods [2, 3]. The vast majority of existing methods [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35] are designed to automate single-step retrosynthetic (SSR) analysis, i.e. predicting a list of compounds from which a target product could be made in one step. A full multi-step route can be determined by iteratively applying SSR methods to the identified precursors until a termination condition is satisfied (e.g., identifying reactants in the database of commercially available compounds). Notably, because each SSR method call creates a list of candidates, iterative application of these methods generates an exponentially growing search space. Graph traversal algorithms such as Monte Carlo Tree Search (MCTS), Depth-First Proof Number (DFPN) search, and A-star (Retro*) have been applied to efficiently traverse this exponential search space [36, 37, 38].
However, evaluation of performance of these SSR methods on full route prediction has been limited by the scarcity of open-source datasets containing valid multi-step routes. The creation of PaRoutes [39], an open-source dataset containing 450k multi-step routes extracted from the United States Patent and Trademark Office (USPTO), marked a major development. Notably, even state-of-the-art SSR models [6, 40] combined with advanced search algorithms (MCTS, DFPN, Retro*) correctly identify multi-step routes (Top-1 accuracy) for only 17% and 10% of the target compounds in n1 and n5, respectively. This performance can be rationalized by recognizing that retrosynthesis is inherently a multi-step problem: the optimal choice of reaction to make a compound depends on subsequent steps in the synthesis. For example, a common pattern in multi-step routes includes: 1) protection of certain functional groups, 2) the desired transformation of unprotected groups, and 3) deprotection of the protected groups. An SSR method applied to the reactant of reaction 2 may output numerous candidate precursors; however, knowing that protective groups are removed in subsequent steps dramatically changes the probability distribution over those precursors.
In this work, we propose a novel approach for direct prediction of multi-step routes, bypassing the need for single-step models and sophisticated exponential graph traversal algorithms. Our DirectMultiStep model demonstrates state-of-the-art performance on n1 and n5 evaluation sets and shows generalizability by successfully predicting routes for FDA-approved drugs absent from the training set. We also discuss the model’s limitations, which may include the need to split longer routes into sub-routes, provide the structure of the starting material, or supply a precursor to the target compound when dealing with transformations not well-represented in the training dataset. Despite these limitations, we believe our results are promising enough to warrant further attention to developing retrosynthetic methods that directly predict multi-step routes.
2 Methods
2.1 Definitions and multi-step route representation
This subsection defines key terms related to our multi-step retrosynthetic methodology. The "target compound" (green in Fig. 1) represents the final product of the multi-step synthesis tree. Starting materials (SM, orange in Fig. 1) are compounds for which no further precursors need to be identified, i.e., they are the leaves of the synthesis tree. The "number of steps" denotes the largest number of reactions from SM to the target compound (tree height). In our approach, routes are represented as recursive dictionaries (Fig. 1) containing the SMILES representation of the molecule and a list of other dictionaries, containing either starting materials or trees leading to precursors of the current node. Removing space and newline characters creates a string representation predicted by our transformer model. We train the DirectMultiStep (DMS) transformer, which takes the target compound, (optionally) starting material, and number of steps as input, and predicts the multi-step route represented as a string.
2.2 Data curation and preparation
A second version of the PaRoutes [39] dataset containing 450k routes is stripped of all metadata and stored as recursive dictionaries representing multi-step routes. The two evaluation sets, n1 and n5, are processed similarly. A training partition is created by removing all permutations (swapping left and right subtrees) of routes in n1 and n5 from the full dataset, resulting in 407k routes. The training dataset is augmented by adding 2 permutations for each route. To train the DMS model which takes starting material as an input, we find all starting materials (leaves) for each tree and store a combination of the target compound with each starting material as a separate entry. As a result, DMS with SM is trained on 3 349 118 inputs, and the DMS without SM on 1 078 951 inputs. All SMILES strings are tokenized by treating each character as a token, and the string representation of the multi-step route is tokenized similarly, treating delimiters of the tree (smiles’, children’, [’, ]’, {’, and }’) as separate tokens. The final vocabulary size is 52 (including start, end, and padding tokens), the largest multi-step route has 1074 tokens, while the largest target compound and starting material have 145 and 135 tokens, respectively.
[…]
Figure 1: The workflow of DirectMultiStep. (a) The SMILES representation of the target compound (green), (optional) starting material (orange), and the number of steps are tokenized, concatenated, and fed into our transformer model. The model predicts a string representation of the multi-step synthesis tree. Spaces are added for clarity, and indentations indicate the levels in the synthesis route (tree). (b) Molecular structures corresponding to the target compound (green), (optional) starting material (orange), and the predicted synthesis tree with structures of all molecules.
Figure 2: Distribution of the relative frequencies of route lengths (in terms of number of steps) in the training dataset before augmentation with permutations (407 148 routes, black), n1 test set (10 000 routes, blue), and n5 test set (10 000 routes, purple). Distribution is split into routes shorter (left subplot) and longer than 6 steps (right subplot).
2.3 Model architecture and training
We present two variants of the DirectMultiStep (DMS) model: DMS-10M and DMS-60M, with approximately 10 million and 60 million parameters, respectively. Both models employ the transformer encoder-decoder architecture [41]. DMS-10M has 6 layers and 8 attention heads in both encoder and decoder, with inputs embedded in 256 dimensions and amplified 3x in the feedforward blocks. DMS-60M has 8 layers and 8 attention heads, with inputs embedded in 512 dimensions and amplified 4x in the feedforward blocks. Both models use GeLU activations [42]. The encoder of DMS with SM receives a concatenation of the tokenized target compound, starting material, and number of steps, while the encoder of DMS without SM receives only the tokenized target compound and number of steps. In all models, the decoder receives output from the encoder and the start token of the multi-step route.
During training, a 10% dropout rate is used in embedding, attention, and feedforward layers. The ADAM optimizer [43] is used with a learning rate scheduler that warms up from 0 to $3\times 10^{-4}$ for DMS-10M and $1\times 10^{-4}$ for DMS-60M during the first 10% of training steps and undergoes cosine decay to $3\times 10^{-5}$ over 20 epochs. Gradients are clipped to 1.0. Finally, during training, each token in the input to the encoders has a 5% chance of being masked. Training requires approximately 200 clock hours using 4 NVIDIA A100-40 GB GPUs on an external cluster. An additional 450 hours were used for experimentation with model architectures and optimization of hyperparameters.
2.4 Beam search for route generation
We utilize beam search with a width of 50 to generate several candidates for the same target compound (and starting material, if provided). The cumulative probability of each beam serves as the ranking metric for prediction accuracy (Top-$K$). After beam search, predicted routes are checked for the validity of all SMILES strings (which are canonicalized), the correctness of the target compound, the presence of starting materials in the stock compound (if applicable), and the absence of repetitions.
2.5 Evaluation metrics and post-processing
Evaluation is performed on all 10 000 routes in both n1 and n5, with the deepest leaf selected as the starting material to be included in the input to the encoder (for models that use starting materials). Top-$K$ accuracy evaluates the percentage of top $K$ predictions precisely matching correct multi-step routes from the test set. Predictions matching any permutation of the correct route are counted as matches.
3 Results and Discussion
3.1 Test accuracy
Top-$K$ accuracy is shown in Table 1 for set-n1 and Table 2 for set-n5. Our DirectMultiStep-SM-10M model achieves state-of-the-art performance, surpassing all previous methods in all Top-$K$ accuracies, with the most significant increase in Top-1 accuracy: 2.2x on n1 and 3.3x on n5. Remarkably, our model’s Top-3 accuracy exceeds the Top-10 accuracy of top tree search methods using single-step retrosynthesis (SSR), highlighting the advantages of the multi-step first approach.
The DirectMultiStep-SM-60M model, despite having six times more parameters than the 10M variant, exhibits lower performance on higher Top-$K$ accuracies, suggesting that the 10M model size is optimal for the current dataset when starting material structure is provided. However, achieving comparable performance without starting material information requires a larger model, as demonstrated by the DirectMultiStep-noSM-60M model’s performance. These findings reveal a trade-off between model generality and size at the current dataset scale: less input information necessitates a larger model to maintain high performance. In all subsequent analysis, we use the routes predicted by the DirectMultiStep-SM-10M model.
Table 1: Top-$K$ accuracy on route test set-n1 (10 000 routes).
Method Top-1 Top-2 Top-3 Top-4 Top-5 Top-10
MCTSa 0.17 0.46 0.49
Retro*a 0.15 0.41 0.45
DFPNa 0.11 0.17 0.17
DirectMultiStep-SM-10M 0.38 0.46 0.50 0.52 0.53 0.56
DirectMultiStep-SM-60M 0.38 0.46 0.48 0.50 0.51 0.52
DirectMultiStep-noSM-60M 0.36 0.42 0.44 0.45 0.45 0.46
aThese data are collected from the 2.0 version of PaRoutes in their GitHub repository [39].
Table 2: Top-$K$ accuracy on route test set-n5 (10 000 routes).
Method Top-1 Top-2 Top-3 Top-4 Top-5 Top-10
MCTSa 0.10 0.28 0.33
Retro*a 0.10 0.30 0.36
DFPNa 0.05 0.07 0.07
DirectMultiStep-SM-10M 0.33 0.40 0.43 0.44 0.46 0.48
DirectMultiStep-SM-60M 0.34 0.41 0.43 0.44 0.45 0.46
DirectMultiStep-noSM-60M 0.32 0.37 0.39 0.40 0.41 0.41
aThese data are collected from the 2.0 version of PaRoutes in their GitHub repository [39]
Importantly, our model requires only one call to obtain multiple routes for each target compound, instead of hundreds of single-step model calls and even more reaction template applications (as shown in PaRoutes [39]). Furthermore, our model’s inference time (beam size 50) on a single GPU is comparable (3-16 seconds depending on the number of steps) to the time required to find the first successful route using previous methods (7-50 seconds [39])
Fig. 3 shows the distribution of Top-1 and Top-10 accuracy over different route lengths on both n1 and n5. Given that 90% of routes in the training partition have 4 or fewer steps Fig. 2), one would expect accuracy to decrease dramatically with increasing route length. However, the performance on routes with 5-8 steps is comparable to that of shorter routes, and the performance on 9-step routes is even comparable to that of 2-step routes. The exceptional accuracy on 9-step routes could be rationalized by the fact that their relative abundance in n1 and n5 does not exceed the relative abundance in the training partition (unlike for routes with length 3-7, as seen in Fig. 3).
Figure 3: Distribution of Top-1 and Top-10 accuracy of predictions with our model on test sets n1 and n5. There is only one route with length 10 in n1, and our model does not predict it correctly. That route is reproduced by splitting it in half, as shown in Fig. 4
[…]
Figure 4: Separation of a 10-step route from set-n1 into two 5-step routes. (a) Correct prediction for the first half of the 10-step route with starting material information. (b) Correct prediction for the second half of the 10-step route with starting material information.
The requirement to provide the starting material structure as input to the model is a limitation, although not as significant as it may seem. Corey’s retrosynthetic framework allows finding at least one route to starting materials by performing one functional group transformation at a time, usually requiring protection of other functional groups. However, such a route will be too long to be efficient in terms of yield (overall yield is the product of yields of individual steps). The art of organic synthesis lies in finding ways to perform several transformations simultaneously using a minimal number of protecting groups. We believe our model provides value as an assistant to this task: finding better ways to get to the target compound from a starting material. The requirement to specify the number of steps as an input becomes a useful option in this context. Moreover, if the target compound is chiral (i.e., it has a mirror image which cannot be superimposed), even during manual retrosynthetic planning, chemists check if chiral centers can be obtained from chiral starting materials (such as naturally occurring amino acids) because developing reactions that create chiral centers in the desired conformation remains an active research area (and subject of 2001 and 2021 Nobel Prizes in Chemistry). An example of such a route is shown in Fig. 6.
3.2 Separation of Long Routes into Shorter Routes
As seen in Fig. 3, our model does not find the single 10-step route in the n1 set. However, this route can be correctly reproduced with rank 1 if split into two 5-step routes and the intermediate compound is provided as the starting material for the first 5 reactions (Fig. 4) and a target compound for the last 5 reactions (Fig. 4). While ideally the necessity to provide an intermediate compound should be avoided, this resembles the bidirectional search technique commonly employed in manual retrosynthetic analysis [44].
3.3 Retrosynthetic planning of pharmaceutical compounds
We demonstrate the generalizability of our model by testing on three FDA-approved drugs: Vonoprazan, Mitapivat, and Daridorexant, which were used for evaluation by Xiong, et al. [45]. These drugs and their intermediates are absent from the training set. Vonoprazan, a potassium competitive acid blocker for Helicobacter pylori infections, was initially proposed with a 2-step route (Fig. 5[46]. However, one of its starting materials is unstable, and the final reaction leads to significant byproducts. Subsequently, an improved 4-step route (Fig. 5) was introduced [47]. Mitapivat, a pyruvate kinase activator for treating hemolytic anemia, has been associated with both a 5-step route (Fig. 5) and a 3-step route (Fig. 5[48, 49]. Daridorexant, an orexin receptor antagonist for adult insomnia, is linked to a 4-step route (Fig. 6[50].
[…]
Figure 5: Literature routes for Vonoprazan and Mitapivat correctly reproduced by our model. Target compounds are in green, starting materials that are given as inputs are colored in orange. Ranks denote the rank of this route when the specified starting material is provided. (a) First literature route for Vonoprazan. The model predicts correctly no matter which starting material is given. (b) Second literature route for Vonoprazan. The model predicts the route correctly only when an immediate precursor to Vonoprazan is given as the target compound. (c) First literature route for Mitapivat. (d) Second literature route for Mitapivat.
Our model correctly reproduces Vonoprazan’s first literature route (Fig. 5) regardless of the starting material provided, with high ranks. However, the model struggles to find the first transformation (counting from the target compound) from the second route. In the first route, the first step is a reductive amination, while in the second route, it’s a reduction of the carbonyl bond in the amide group. Reductive amination is much more common for creating C-N bonds, so the training set is heavily biased toward it. As a result, all routes predicted by our model start as in Fig. 5. However, if we provide the precursor to Vonoprazan in the second route as the target compound, the model correctly reproduces the full route (Fig. 5). This highlights an important limitation of our model: it may not perform as well on routes involving transformations that are inadequately represented in the training dataset. We envision that as larger multi-step datasets become available, this problem will naturally disappear.
Figure 6: Daridorexant literature multi-step routes and predictions from the model with starting material information. The compounds that are missing from the prediction are in red.
Our model correctly reproduces both routes for Mitapivat (Fig. 5 and Fig. 5). In contrast, the route for Daridorexant (Fig. 6) is only partially predicted: our model does not split the right child of the root node (Daridorexant itself) into the two compounds highlighted in red. All other intermediates and starting materials are predicted correctly. The reason for this behavior is not well understood, but it can be rationalized by looking at the distribution of the number of leaves at root nodes: 73% of root nodes (target compounds) have at least 1 leaf (see SI Fig. 7). Such dependence on the distributions of the training set comprises an important limitation of the multi-step first approach. This issue would not occur with methods based on single-step prediction because the tree search framework would not halt at compounds outside the stock compound set before reaching the maximum iterations. Ensuring that multi-step first models reproduce that aspect of SSR methods constitutes a direction for future research.
4 Conclusion
In this work, we introduced DirectMultiStep, a transformer-based model that directly generates multi-step synthetic routes for a given target compound. By predicting the entire route as a single string, our approach bypasses the need for iterative single-step predictions and exponential search space traversal. The model accommodates specific conditions such as the desired number of steps and starting materials, enabling efficient route planning.
DirectMultiStep outperforms state-of-the-art methods on the PaRoutes dataset, achieving a 2.2x improvement in Top-1 accuracy on the n1 test set and a 3.3x improvement on the n5 test set. The model also successfully predicts routes for FDA-approved drugs not included in the training data, showcasing its generalization capabilities. However, the model’s performance may be impacted by the suboptimal diversity of the current training set, particularly for less common reaction types.
Despite these limitations, the DirectMultiStep approach presents a promising direction towards fully automated retrosynthetic planning. As larger and more diverse multi-step datasets become available, the model’s performance is expected to improve further. Future research should focus on addressing the limitations identified, such as the need to split longer routes into sub-routes and the dependence on the distributions of the training set. Overall, this work demonstrates the potential of transformer-based models for efficient and scalable computer-aided synthesis planning.
5 Broad Impact
Chemical synthesis of target compounds of interest forms the routine for millions of organic chemists around the world, including graduate students in PhD programs and synthetic chemists in pharmaceutical companies. The ability to find shorter (and thus more efficient) routes to any desired compound is an enticing challenge with the potential of saving on labor costs and chemical waste. As a result, multi-step prediction methods have the potential to improve the lives of millions of people, and potentially contribute to faster advances in the development of new drugs and materials. We hope that our work constitutes an important advancement in that direction.
Quite unfortunately, however, there is no fundamental chemical difference between how safe and useful compounds (such as drugs) and malign compounds (such as chemical weapons) are synthesized. Both require application of the same set of rules of organic synthesis, and so the ability to make drugs faster and more efficiently inherently brings along the ability to make poisons faster and more efficiently. One cannot impose restrictions on the tokens (chemical elements) learned by the model, as even some combinations of nitrogen and oxygen (which are widespread in proteins and DNA) may be cytotoxic. Therefore, unless the humanity at large decides that the benefits of automated retrosynthesis do not outweigh its potential harms, the most we can do is communicate the dangers clearly and perhaps impose restrictions that would still require a trained (which often accompanies training in ethics) chemist to use our model. Besides, all meta-data of reaction conditions and catalysts are removed from the predictions of our model, which would present a deliberate barrier against malicious uses.
6 Code and Data Availability
The reaction route dataset is from the 2.0 version of PaRoutes in their GitHub repository [39], available under the Apache License 2.0. Code to process dataset, implementation of model architecture, code for training, generation, and evaluation are available under MIT License at https://github.com/batistagroup/DirectMultiStep.
7 Acknowledgments
The authors acknowledge a generous allocation of high-performance computing time from NERSC. The development of the methodology was supported by the NSF CCI grant (VSB, Award Number 2124511).

### v2

@@
-DirectMultiStep: Direct Route Generation for Multi-Step Retrosynthesis
+DirectMultiStep: Direct Route Generation for Multistep Retrosynthesis
 Abstract
 Traditional computer-aided synthesis planning (CASP) methods rely on iterative single-step predictions, leading to exponential search space growth that limits efficiency and scalability.
-We introduce a transformer-based model that directly generates multi-step synthetic routes as a single string by conditionally predicting each molecule based on all preceding ones.
-The model accommodates specific conditions such as the desired number of steps and starting materials, outperforming state-of-the-art methods on the PaRoutes dataset with a 2.2x improvement in Top-1 accuracy on the n1 test set and a 3.3x improvement on the n5 test set.
+We introduce a series of transformer-based models, utilizing mixture of experts approach, that directly generate multistep synthetic routes as a single string by conditionally predicting each molecule based on all preceding ones.
+Our models can accommodate specific conditions such as the desired number of steps and starting materials, with the top-performing DMS-Flex (Duo) surpassing state-of-the-art methods on the PaRoutes dataset with a 2.5x improvement in Top-1 accuracy on the n1 test set and a 3.9x improvement on the n5 test set.
 It also successfully predicts routes for FDA-approved drugs not included in the training data, showcasing its generalization capabilities.
-While the current suboptimal diversity of the training set may impact performance on less common reaction types, our approach presents a promising direction towards fully automated retrosynthetic planning.
+While the current suboptimal diversity of the training set may impact performance on less common reaction types, our multistep-first approach presents a promising direction towards fully automated retrosynthetic planning.
 1 Introduction
@@
 It’s worth noting that many commercially available compounds can be synthesized from other commercially available compounds, adding another layer of complexity to the decision-making process.
-A significant contribution to the development of algorithmic frameworks for identifying synthetic routes leading to any desired compound (subject of the 1990 Nobel Prize in Chemistry) was made by Elias James Corey.
-Corey’s framework, now known as retrosynthetic analysis, begins with identifying atoms that would serve as reaction centers.
+Algorithmic frameworks for identifying synthetic routes were envisioned by Vleduts [1] and further formalized by Elias James Corey (subject of the 1990 Nobel Prize in Chemistry) into what is now known as retrosynthetic analysis.
+This framework begins with identifying atoms that would serve as reaction centers.
 Disconnecting bonds between these centers results in the formation of hypothetical fragments (called synthons) from which a precursor molecule can be created.
-This mapping from fragments to actual molecules is one-to-many because there is usually more than one functional group that could participate in any given type of reaction.
+This mapping from synthons to actual molecules is one-to-many because there is usually more than one functional group that could participate in any given type of reaction.
 Importantly, meticulous application of Corey’s framework (i.e., systematically breaking small subsets of bonds) will eventually lead to commercially available starting materials.
@@
 predicting a list of compounds from which a target product could be made in one step.
-A full multi-step route can be determined by iteratively applying SSR methods to the identified precursors until a termination condition is satisfied (e.g., identifying reactants in the database of commercially available compounds).
+A full multistep route can be determined by iteratively applying SSR methods to the identified precursors until a termination condition is satisfied (e.g., identifying reactants in the database of commercially available compounds).
 Notably, because each SSR method call creates a list of candidates, iterative application of these methods generates an exponentially growing search space.
 Graph traversal algorithms such as Monte Carlo Tree Search (MCTS), Depth-First Proof Number (DFPN) search, and A-star (Retro*) have been applied to efficiently traverse this exponential search space [37, 38, 39].
-However, evaluation of performance of these SSR methods on full route prediction has been limited by the scarcity of open-source datasets containing valid multi-step routes.
-The creation of PaRoutes [39], an open-source dataset containing 450k multi-step routes extracted from the United States Patent and Trademark Office (USPTO), marked a major development.
-Notably, even state-of-the-art SSR models [6, 40] combined with advanced search algorithms (MCTS, DFPN, Retro*) correctly identify multi-step routes (Top-1 accuracy) for only 17% and 10% of the target compounds in n1 and n5, respectively.
-This performance can be rationalized by recognizing that retrosynthesis is inherently a multi-step problem: the optimal choice of reaction to make a compound depends on subsequent steps in the synthesis.
-For example, a common pattern in multi-step routes includes: 1) protection of certain functional groups, 2) the desired transformation of unprotected groups, and 3) deprotection of the protected groups.
-An SSR method applied to the reactant of reaction 2 may output numerous candidate precursors; however, knowing that protective groups are removed in subsequent steps dramatically changes the probability distribution over those precursors.
-In this work, we propose a novel approach for direct prediction of multi-step routes, bypassing the need for single-step models and sophisticated exponential graph traversal algorithms.
-Our DirectMultiStep model demonstrates state-of-the-art performance on n1 and n5 evaluation sets and shows generalizability by successfully predicting routes for FDA-approved drugs absent from the training set.
-We also discuss the model’s limitations, which may include the need to split longer routes into sub-routes, provide the structure of the starting material, or supply a precursor to the target compound when dealing with transformations not well-represented in the training dataset.
-Despite these limitations, we believe our results are promising enough to warrant further attention to developing retrosynthetic methods that directly predict multi-step routes.
-2 Methods
-2.1 Definitions and multi-step route representation
-This subsection defines key terms related to our multi-step retrosynthetic methodology.
-The "target compound" (green in Fig.
-1) represents the final product of the multi-step synthesis tree.
-Starting materials (SM, orange in Fig.
-1) are compounds for which no further precursors need to be identified, i.e., they are the leaves of the synthesis tree.
-The "number of steps" denotes the largest number of reactions from SM to the target compound (tree height).
-In our approach, routes are represented as recursive dictionaries (Fig.
+More sophisticated search designs include utilization of hyper-graph [40], graph-based neural networks [41], reinforcement learning [42], or incorporation of previous reactions as context for single-step predictions [43].
+Evaluation of performance of these SSR methods on full route prediction has been limited by the scarcity of open-source datasets containing valid multistep routes.
+The creation of PaRoutes [44], an open-source dataset containing 450k multistep routes (163k excluding duplicates and permutations, along with two test sets of 10k routes each) extracted from the United States Patent and Trademark Office (USPTO), marked a major development.
+Notably, even state-of-the-art SSR models [7, 45] combined with advanced search algorithms (MCTS, DFPN, Retro*) correctly identify multistep routes (Top-1 accuracy) for only 17% and 10% of the target compounds in n1 and n5 test sets, respectively.
+Such performance could be rationalized by recognizing that retrosynthesis is inherently a multistep problem: the optimal choice of reaction to make a compound depends on subsequent steps in the synthesis.
+For example, a common pattern in multistep routes includes: (1) protection of certain functional groups, (2) the desired transformation of unprotected groups, and (3) deprotection of the protected groups.
+An SSR method applied to the reactant of reaction (2) may output numerous candidate precursors; however, knowing that protective groups are removed in subsequent steps dramatically changes the probability distribution over those precursors.
+In this work, we propose a novel approach for direct prediction of multistep routes, bypassing the need for single-step models and sophisticated exponential graph traversal algorithms.
+We begin with a brief overview of the methodology and evaluation metrics.
+We proceed by demonstrating and discussing performance of DMS models on diverse n1 and n5 evaluation sets (10 000 routes each).
+We demonstrate further generalizability by successfully predicting experimental synthetic routes for FDA-approved drugs.
+We conclude by discussing the limitations of our approach.
+A more elaborate description of model architectures and rationale for certain design choices is given at the end of the paper.
+2 Methodology Overview
+2.1 Definitions
+Target compound (blue in Fig.
+1) is the final (desired) product of the multistep synthesis tree.
+Starting material, SM (red in Fig.
+1) is a compound for which no further precursors need to be identified, i.e., a leaf of the synthesis tree.
+The number of steps is the largest number of reactions from SM to the target compound (tree height).
+2.2 Route Representation
+Routes are represented as recursive dictionaries (Fig.
 1) containing the SMILES representation of the molecule and a list of other dictionaries, containing either starting materials or trees leading to precursors of the current node.
-Removing space and newline characters creates a string representation predicted by our transformer model.
-We train the DirectMultiStep (DMS) transformer, which takes the target compound, (optionally) starting material, and number of steps as input, and predicts the multi-step route represented as a string.
-2.2 Data curation and preparation
-A second version of the PaRoutes [39] dataset containing 450k routes is stripped of all metadata and stored as recursive dictionaries representing multi-step routes.
-The two evaluation sets, n1 and n5, are processed similarly.
-A training partition is created by removing all permutations (swapping left and right subtrees) of routes in n1 and n5 from the full dataset, resulting in 407k routes.
-The training dataset is augmented by adding 2 permutations for each route.
-To train the DMS model which takes starting material as an input, we find all starting materials (leaves) for each tree and store a combination of the target compound with each starting material as a separate entry.
-As a result, DMS with SM is trained on 3 349 118 inputs, and the DMS without SM on 1 078 951 inputs.
-All SMILES strings are tokenized by treating each character as a token, and the string representation of the multi-step route is tokenized similarly, treating delimiters of the tree (smiles’, children’, [’, ]’, {’, and }’) as separate tokens.
-The final vocabulary size is 52 (including start, end, and padding tokens), the largest multi-step route has 1074 tokens, while the largest target compound and starting material have 145 and 135 tokens, respectively.
+Removing space and newline characters creates a string representation.
+Our models predict that string by taking SMILES of the target compound, (optionally) starting material, and (optionally) number of steps as input.
 […]
@@
 Figure 1: The workflow of DirectMultiStep.
-(a) The SMILES representation of the target compound (green), (optional) starting material (orange), and the number of steps are tokenized, concatenated, and fed into our transformer model.
-The model predicts a string representation of the multi-step synthesis tree.
+(a) The SMILES representation of the target compound (blue), (optional) starting material (red), and the number of steps are tokenized, concatenated, and fed into our transformer model.
+The model predicts a string representation of the multistep synthesis tree.
 Spaces are added for clarity, and indentations indicate the levels in the synthesis route (tree).
-(b) Molecular structures corresponding to the target compound (green), (optional) starting material (orange), and the predicted synthesis tree with structures of all molecules.
-Figure 2: Distribution of the relative frequencies of route lengths (in terms of number of steps) in the training dataset before augmentation with permutations (407 148 routes, black), n1 test set (10 000 routes, blue), and n5 test set (10 000 routes, purple).
+(b) Molecular structures corresponding to the target compound (blue), (optional) starting material (red), and the predicted synthesis tree with structures of all molecules.
+2.3 Evaluation Metrics and Post-Processing
+Evaluation is performed on n1 and n5 (10 000 routes each) evaluation sets from PaRoutes [44], which have been created by selecting the most diverse routes from the USPTO.
+Fig.
+2 shows the distribution of route lengths in training set, n1, and n5.
+The distribution in both n1 and n5 is shifted towards longer routes; for example, the fraction of routes with 4 steps or more is 15% for training set, 27% for n1 and 51% for n5.
+Figure 2: Distribution of the relative frequencies of route lengths (in terms of number of steps) in the training dataset before augmentation with permutations (163 689 routes, black), n1 test set (10 000 routes, blue), and n5 test set (10 000 routes, purple).
 Distribution is split into routes shorter (left subplot) and longer than 6 steps (right subplot).
-2.3 Model architecture and training
-We present two variants of the DirectMultiStep (DMS) model: DMS-10M and DMS-60M, with approximately 10 million and 60 million parameters, respectively.
-Both models employ the transformer encoder-decoder architecture [41].
-DMS-10M has 6 layers and 8 attention heads in both encoder and decoder, with inputs embedded in 256 dimensions and amplified 3x in the feedforward blocks.
-DMS-60M has 8 layers and 8 attention heads, with inputs embedded in 512 dimensions and amplified 4x in the feedforward blocks.
-Both models use GeLU activations [42].
-The encoder of DMS with SM receives a concatenation of the tokenized target compound, starting material, and number of steps, while the encoder of DMS without SM receives only the tokenized target compound and number of steps.
-In all models, the decoder receives output from the encoder and the start token of the multi-step route.
-During training, a 10% dropout rate is used in embedding, attention, and feedforward layers.
-The ADAM optimizer [43] is used with a learning rate scheduler that warms up from 0 to $3\times 10^{-4}$ for DMS-10M and $1\times 10^{-4}$ for DMS-60M during the first 10% of training steps and undergoes cosine decay to $3\times 10^{-5}$ over 20 epochs.
-Gradients are clipped to 1.0.
-Finally, during training, each token in the input to the encoders has a 5% chance of being masked.
-Training requires approximately 200 clock hours using 4 NVIDIA A100-40 GB GPUs on an external cluster.
-An additional 450 hours were used for experimentation with model architectures and optimization of hyperparameters.
-2.4 Beam search for route generation
-We utilize beam search with a width of 50 to generate several candidates for the same target compound (and starting material, if provided).
-The cumulative probability of each beam serves as the ranking metric for prediction accuracy (Top-$K$).
-After beam search, predicted routes are checked for the validity of all SMILES strings (which are canonicalized), the correctness of the target compound, the presence of starting materials in the stock compound (if applicable), and the absence of repetitions.
-2.5 Evaluation metrics and post-processing
-Evaluation is performed on all 10 000 routes in both n1 and n5, with the deepest leaf selected as the starting material to be included in the input to the encoder (for models that use starting materials).
-Top-$K$ accuracy evaluates the percentage of top $K$ predictions precisely matching correct multi-step routes from the test set.
-Predictions matching any permutation of the correct route are counted as matches.
+The structure of target compound, starting material (optionally, the deepest leaf of the correct route is chosen during evaluation), and number of steps is passed to the transformer encoder.
+A beam search with a width of 50 is employed to generate 50 candidate routes by sampling next tokens from the decoder.
+Within beam search, during the prediction of each token the cumulative log likelihoods are normalized by the sequence length at the time the new token is sampled.
+This adjustment ensures that shorter routes with simpler compounds are not disproportionately favored while maintaining a balance between route length and overall sequence probability.
+After beam search finds end-of-sequence tokens for all 50 candidates, predicted routes are checked for validity of all SMILES strings (which are canonicalized), the presence of starting materials in the stock set (following procedure of PaRoutes taken to be the set of all leaves in n1 or n5, and absence of repetitions (which can be present as permutations of other routes).
+Remaining routes are compared against the true (experimental) route.
+The reproduction of experimental routes is presented as Top-$K$ accuracy, i.e.
+presence of correct route ranked $K$ or lower.
+Importantly, the Top-$K$ accuracy is evaluated against all 10 000 targets, not just the ones for which routes are found.
+2.4 Models Presented
+As a baseline, we train a conventional transformer-based sequence-to-sequence model with 6 encoder and 6 decoder layers (DMS-Flash) with structures of starting materials provided during training.
+To achieve comparable performance without starting material as a hint, a significantly larger model with 12 encoder and 36 decoder layers is needed (DMS-Deep).
+We also present the first, to our knowledge, application of mixture-of-experts (MoE [46, 47, 48]) architecture to a chemical problem.
+To create MoE model, one replaces the 2-layer MLP that follows every attention block with an ensemble of 2-layer MLPs (called experts).
+A preceding router determines (for each input) the $n$-most relevant (active) experts, and passes the input through each of them.
+Output of each active expert is added with weights determined by the router.
+Notably, the number of experts active during inference does not have to be equal to the number of experts active during training, although the latter tends to be optimal.
+We present two MoE models, the DMS-Flex (trained with SM) and DMS-Wide (trained without SM).
+The relevant architectural details of models are given in Table 1.
+Table 1: Specifications of architecture for different model variants.
+DMS Model Encoder Decoder Hidden FF Total Active Experts
+Variant Layers Layers Dimension Multipliera Experts Training Generation
+Flash 6 6 256 3 1 1 1
+Explorer 6 6 256 3 3 2 2
+Explorer XL 6 24 256 3 3 2 2
+Flex (Mono) 6 6 256 3 3 2 1
+Flex (Duo) 6 6 256 3 3 2 2
+Deep 12 36 256 3 1 1 1
+Wide 12 12 256 3 3 2 2
+a an integer by which the hidden dimension is multiplied in feed-forward layers after attention.
 3 Results and Discussion
-3.1 Test accuracy
+3.1 Top-K Accuracy on Evaluation Sets
+3.1.1 Vanilla Transformers
 Top-$K$ accuracy is shown in Table 2 for set-n1 and Table 3 for set-n5.
-Our DirectMultiStep-SM-10M model achieves state-of-the-art performance, surpassing all previous methods in all Top-$K$ accuracies, with the most significant increase in Top-1 accuracy: 2.2x on n1 and 3.3x on n5.
-Remarkably, our model’s Top-3 accuracy exceeds the Top-10 accuracy of top tree search methods using single-step retrosynthesis (SSR), highlighting the advantages of the multi-step first approach.
-The DirectMultiStep-SM-60M model, despite having six times more parameters than the 10M variant, exhibits lower performance on higher Top-$K$ accuracies, suggesting that the 10M model size is optimal for the current dataset when starting material structure is provided.
-However, achieving comparable performance without starting material information requires a larger model, as demonstrated by the DirectMultiStep-noSM-60M model’s performance.
-These findings reveal a trade-off between model generality and size at the current dataset scale: less input information necessitates a larger model to maintain high performance.
-In all subsequent analysis, we use the routes predicted by the DirectMultiStep-SM-10M model.
+The smallest DMS-Flash model (which receives structure of SM during generation) achieves state-of-the-art performance, surpassing all previous methods in all Top-$K$ accuracies, with the most significant increase in Top-1 accuracy: 2.4x on n1 and 3.7x on n5.
+To achieve comparable performance without providing the structure of the SM, a significantly larger model DMS-Deep is needed.
+This establishes a trade-off between size of the model (and thus generation speed) and the amount of information it receives.
+The limitations of the necessity to provide the structure of the SM are discussed in Section 3.1.3.
+Quite remarkably, a Flash model trained with structures of SM may generate routes even when SMs are not provided (Tables 2-3).
+While that results in a 20% drop in Top-1 accuracy, it is significantly better than a model with the same architecture that is trained without SM (Tables S6-S7).
+In other words, providing structure of a correct starting material helps the model learn better and improves performance even if the starting material is not used during evaluation.
 Table 2: Top-$K$ accuracy on route test set-n1 (10 000 routes).
-Method Top-1 Top-2 Top-3 Top-4 Top-5 Top-10
+Method Top-1 Top-5 Top-10
 MCTSa 0.17 0.46 0.49
@@
 DFPNa 0.11 0.17 0.17
-DirectMultiStep-SM-10M 0.38 0.46 0.50 0.52 0.53 0.56
-DirectMultiStep-SM-60M 0.38 0.46 0.48 0.50 0.51 0.52
-DirectMultiStep-noSM-60M 0.36 0.42 0.44 0.45 0.45 0.46
+DMS SM SM
+Models Size Steps (Train) (Gen.) Top-1 Top-2 Top-3 Top-4 Top-5 Top-10
+Flash 10M 0.41 0.50 0.54 0.56 0.58 0.60
+Flash 10M 0.32 0.39 0.42 0.44 0.45 0.47
+Explorer 19M 0.36 0.45 0.48 0.51 0.52 0.55
+Explorer 19M 0.27 0.33 0.36 0.37 0.39 0.40
+Explorer-XL 50M 0.32 0.40 0.44 0.45 0.46 0.48
+Flex (Mono) 19M 0.32 0.40 0.43 0.44 0.45 0.48
+Flex (Duo) 19M 0.43 0.52 0.55 0.57 0.58 0.61
+Flex (Duo) 19M 0.34 0.41 0.44 0.45 0.46 0.47
+Deep 41M 0.36 0.44 0.47 0.49 0.50 0.52
+Wide 38M 0.38 0.46 0.49 0.50 0.51 0.53
 aThese data are collected from the 2.0 version of PaRoutes in their GitHub repository [44].
 Table 3: Top-$K$ accuracy on route test set-n5 (10 000 routes).
-Method Top-1 Top-2 Top-3 Top-4 Top-5 Top-10
+Method Top-1 Top-5 Top-10
 MCTSa 0.10 0.28 0.33
@@
 DFPNa 0.05 0.07 0.07
-DirectMultiStep-SM-10M 0.33 0.40 0.43 0.44 0.46 0.48
-DirectMultiStep-SM-60M 0.34 0.41 0.43 0.44 0.45 0.46
-DirectMultiStep-noSM-60M 0.32 0.37 0.39 0.40 0.41 0.41
-aThese data are collected from the 2.0 version of PaRoutes in their GitHub repository [39]
-Importantly, our model requires only one call to obtain multiple routes for each target compound, instead of hundreds of single-step model calls and even more reaction template applications (as shown in PaRoutes [39]).
-Furthermore, our model’s inference time (beam size 50) on a single GPU is comparable (3-16 seconds depending on the number of steps) to the time required to find the first successful route using previous methods (7-50 seconds [39])
+DMS SM SM
+Models Size Steps (Train) (Gen.) Top-1 Top-2 Top-3 Top-4 Top-5 Top-10
+Flash 10M 0.37 0.44 0.47 0.49 0.51 0.53
+Flash 10M 0.30 0.36 0.39 0.40 0.41 0.43
+Explorer 19M 0.33 0.40 0.43 0.45 0.46 0.49
+Explorer 19M 0.25 0.31 0.34 0.35 0.36 0.37
+Explorer-XL 50M 0.31 0.38 0.41 0.42 0.43 0.45
+Flex (Mono) 19M 0.28 0.34 0.36 0.38 0.39 0.41
+Flex (Duo) 19M 0.39 0.46 0.49 0.51 0.52 0.54
+Flex (Duo) 19M 0.32 0.38 0.40 0.42 0.42 0.44
+Deep 41M 0.33 0.40 0.43 0.45 0.46 0.48
+Wide 38M 0.35 0.42 0.44 0.46 0.47 0.49
+aThese data are collected from the 2.0 version of PaRoutes in their GitHub repository [44].
+Remarkably, the Top-2 accuracy of DMS-Flash and the Top-4 accuracy of DMS-Deep exceed the Top-10 accuracy of all search methods that utilize iterative application of single-step models, highlighting the advantages of the multistep-first approach.
+To understand this performance difference, it helps to notice that iterative application of single-step calls results in a breadth-first search-like exploration of a node, as all reactants leading to a compound are generated at the same time.
+Our approach, however, is a depth-first search-like: the whole left sub-tree of target compound in Fig.
+1 is predicted before the immediate right child.
+This provides an extra context for the prediction of the second precursor to the target compound (or any other right child), which reduces the search space and finds more relevant compounds faster.
+A few words on the relevance of Top-$K$ accuracy as the primary evaluation metric are in order.
+One might question whether the number of targets for which valid (i.e., having valid SMILES) routes are found (or solvability rate if all SMs are in a given stock set) and the average length of routes predicted should be considered.
+It is essential to realize that even if you ensure that every single-step transformation corresponds to a known reaction by template matching, the route may still be experimentally not feasible simply because atoms that are unchanged (i.e., the ones ignored by the template) contribute to the energies of the reacting orbitals through inductive and resonance effects, thereby changing the energetic favorability of the reaction.
+As a result, short of performing highly-accurate ab initio calculations of energies (after corresponding geometry optimization) with all relevant solvation effects and considering thermal corrections, there is no computational metric that would assess experimental feasibility of the predicted route.
+In addition, one could argue that the ability to discover an experimentally verified route is a prerequisite for finding shorter, but still feasible, routes.
+Therefore, Top-$K$ accuracy is the closest proxy to route quality.
+Additionally, the ability to predict correct routes with better ranks is a significant improvement because it reduces the number of candidates to be passed through the aforementioned highly expensive energy calculations.
+Finally, one could argue that single-step methods do not provide inherent ranking of candidates, so the performance of search methods may depend on the choice of ranking metric.
+However, that is yet another benefit of our approach: there is no need in devising another algorithm or model for route ranking, the beam probabilities are sufficiently accurate.
+The superior performance of DMS models does not incur additional computational cost.
+A search with a beam size of 50 takes 3-16 seconds (depending on the number of steps) with a DMS-Flash or a DMS-Flex model on a single GPU (NVIDIA M4000), which is comparable to the time required to find the first successful route using previous methods (7-50 seconds [44]).
+Figure 3: Distribution of Top-1 and Top-10 accuracy of predictions with DMS-Flex (Duo) on test sets n1 and n5.
+There is only one route with length 10 in n1, and DMS-Flex (Duo) does not predict it correctly.
+That route is reproduced by splitting it in half, as shown in Fig.
+4
+3.1.2 MoE Transformers
+The DMS-Flex (Flexible), a mixture-of-experts model trained with two out of three experts active, provides the largest performance improvement when evaluated with two experts active (Duo mode): 2.5x increase in Top-1 on n1 and 3.9x on n5.
+Evaluating DMS-Flex with one expert active (Mono) mode worsens the performance by 30%; given that the number of valid routes (Tables S1-S2) does not change significantly, one could consider the Mono mode as a source of diversity in predictions.
+Notably, even though the DMS-Flex is trained with SM, it can be evaluated without providing SM (empty padding tokens are fed into the encoder instead) which still results in 34% Top-1 accuracy and 47+Top-10.
+Evaluating DMS-Flex without providing the number of steps as an input drops performance significantly, but a different model with the same architecture trained without number of steps as an input, DMS-Explorer (Fig.
+S1 shows the distribution of route lengths predicted by Explorer relative to the length of an experimental route), results in only a 8% drop in accuracy.
+Using a larger DMS-Wide model trained without SM provides comparable performance to that of DMS-Deep.
+An important advantage of DMS-Wide is that tokens passed through different active experts can be processed on a GPU in parallel, which enables faster predictions.
+In addition, DMS-Wide has fewer layers, so fewer consecutive calculations have to be performed; in practice, doubling the number of encoder layers results in doubling of generation time.
+One might also conclude that similar performance of DMS-Deep and DMS-Wide, which have comparable parameter counts, implies that the expressive power mostly depends on the number of parameters, and not on the extra non-linearities provided by larger number of layers.
+Our experimentation with the increase in the number of experts, the number of encoder/decoder layers, or the feed-forward multiplier evaluated on small subsets of n1 or n5, show no significant improvement in performance (Tables S4-S7), which most likely means that the 20 M (with SM) or 40 M (without SM) are optimal given the current size of the training set.
 Fig.
 3 shows the distribution of Top-1 and Top-10 accuracy over different route lengths on both n1 and n5.
-Given that 90% of routes in the training partition have 4 or fewer steps Fig.
-2), one would expect accuracy to decrease dramatically with increasing route length.
+Given that 85% of routes in the training partition have 4 or fewer steps Fig.
+2), one would expect the accuracy to decrease dramatically with increasing route length.
 However, the performance on routes with 5-8 steps is comparable to that of shorter routes, and the performance on 9-step routes is even comparable to that of 2-step routes.
@@
 3).
-Figure 3: Distribution of Top-1 and Top-10 accuracy of predictions with our model on test sets n1 and n5.
-There is only one route with length 10 in n1, and our model does not predict it correctly.
-That route is reproduced by splitting it in half, as shown in Fig.
-4
+Top-1 and Top-10 accuracies of other DMS models for different route lengths are given in Tables S8-S11.
 […]
@@
 Figure 4: Separation of a 10-step route from set-n1 into two 5-step routes.
-(a) Correct prediction for the first half of the 10-step route with starting material information.
-(b) Correct prediction for the second half of the 10-step route with starting material information.
-The requirement to provide the starting material structure as input to the model is a limitation, although not as significant as it may seem.
-Corey’s retrosynthetic framework allows finding at least one route to starting materials by performing one functional group transformation at a time, usually requiring protection of other functional groups.
-However, such a route will be too long to be efficient in terms of yield (overall yield is the product of yields of individual steps).
+(a) Correct prediction from DMS-Flex (Duo) for the first half of the 10-step route with starting material information (red).
+(b) Correct prediction from DMS-Flex (Duo) for the second half of the 10-step route with starting material information.
+3.1.3 Use Cases for Different DMS Model Variants
+The requirement to provide the starting material structure as input to the model might, in theory, seem as a limitation, limiting the relevance of DMS-Flash/Flex models.
+However, in practice, a chemist often can know the structure of a desired starting material because Corey’s retrosynthetic framework allows finding at least one route to starting materials by performing one functional group transformation at a time.
+Such deterministically found routes will often be too long to be efficient in terms of yield (overall yield is the product of yields of individual steps).
 The art of organic synthesis lies in finding ways to perform several transformations simultaneously using a minimal number of protecting groups.
-We believe our model provides value as an assistant to this task: finding better ways to get to the target compound from a starting material.
+As such, we believe our SM-based models provide value as an assistant to this task: finding better ways to get to the target compound from a starting material.
 The requirement to specify the number of steps as an input becomes a useful option in this context.
@@
 An example of such a route is shown in Fig.
-6.
+7.
+In addition, one can use more computationally expensive DMS-Wide/Deep models to discover potential starting materials, and then use those structures as inputs to DMS-Flash/Flex to generate suggestions for each starting material of interest.
+To lift the necessity to specify desired route length, a user also has a choice between running Flex/Wide models several times with different desired route lengths or simply running Explorer-XL once.
+Of course, ideally one would prefer a single model that predicts the whole route with just from the structure of the target compound, without number of steps or starting material provided as an extra input.
+We believe better performance could be attainable in the future once bigger datasets become available.
+Finding ways to improve performance even with current dataset constitutes an important area of further research.
 3.2 Separation of Long Routes into Shorter Routes
 As seen in Fig.
-3, our model does not find the single 10-step route in the n1 set.
+3, DMS-Flex (Duo) does not find the single 10-step route in the n1 set.
 However, this route can be correctly reproduced with rank 1 if split into two 5-step routes and the intermediate compound is provided as the starting material for the first 5 reactions (Fig.
@@
 While ideally the necessity to provide an intermediate compound should be avoided, this resembles the bidirectional search technique commonly employed in manual retrosynthetic analysis [49].
+In contrast, models trained without SM, such as DMS-Deep and DMS-Wide, can replicate the second 5-step sequence in Fig.
+4 but not the first (Fig.
+4).
+[…]
+[…]
+[…]
+[…]
+Figure 5: Literature routes for Vonoprazan and Mitapivat correctly reproduced by DMS-Flex (Duo).
+Target compounds are in blue, starting materials that are given as inputs are colored in red.
+Ranks denote the rank of this route when the specified starting material is provided.
+(a) First literature route for Vonoprazan.
+The model predicts correctly no matter which starting material is given.
+(b) Second literature route for Vonoprazan.
+The model predicts the route correctly only when an immediate precursor to Vonoprazan is given as the target compound.
+(c) First literature route for Mitapivat.
+(d) Second literature route for Mitapivat.
 3.3 Retrosynthetic planning of pharmaceutical compounds
-We demonstrate the generalizability of our model by testing on three FDA-approved drugs: Vonoprazan, Mitapivat, and Daridorexant, which were used for evaluation by Xiong, et al.
+For another generalizability assessment of our model, we test it on three FDA-approved drugs: Vonoprazan, Mitapivat, and Daridorexant, which were used for evaluation by Xiong, et al.
 [50].
@@
 Daridorexant, an orexin receptor antagonist for adult insomnia, is linked to a 4-step route (Fig.
-6[50].
-[…]
-[…]
-[…]
-[…]
-Figure 5: Literature routes for Vonoprazan and Mitapivat correctly reproduced by our model.
-Target compounds are in green, starting materials that are given as inputs are colored in orange.
-Ranks denote the rank of this route when the specified starting material is provided.
-(a) First literature route for Vonoprazan.
-The model predicts correctly no matter which starting material is given.
-(b) Second literature route for Vonoprazan.
-The model predicts the route correctly only when an immediate precursor to Vonoprazan is given as the target compound.
-(c) First literature route for Mitapivat.
-(d) Second literature route for Mitapivat.
-Our model correctly reproduces Vonoprazan’s first literature route (Fig.
+7[55].
+Both vanilla transformers (DMS-Flash) and MoE models (DMS-Flex Duo) correctly reproduce Vonoprazan’s first literature route (Fig.
 5) regardless of the starting material provided, with high ranks.
-However, the model struggles to find the first transformation (counting from the target compound) from the second route.
-In the first route, the first step is a reductive amination, while in the second route, it’s a reduction of the carbonyl bond in the amide group.
-Reductive amination is much more common for creating C-N bonds, so the training set is heavily biased toward it.
-As a result, all routes predicted by our model start as in Fig.
+However, DMS-Flash struggles to find the first transformation (counting from the target compound) in the second route.
+This inability could be attributed to the fact that the reduction of amide carbonyl (the step leading to Vonoprazan in second route) is less frequently represented in the PaRoutes dataset than the much more common reductive amination (the step leading to Vonoprazan in first route).
+As a result, most routes predicted by DMS-Flash start with a reductive amination as in Fig.
 5.
-However, if we provide the precursor to Vonoprazan in the second route as the target compound, the model correctly reproduces the full route (Fig.
+However, if we provide the precursor to Vonoprazan in the second route as the target compound, the model correctly reproduces the remaining route (Fig.
 5).
-This highlights an important limitation of our model: it may not perform as well on routes involving transformations that are inadequately represented in the training dataset.
-We envision that as larger multi-step datasets become available, this problem will naturally disappear.
-Figure 6: Daridorexant literature multi-step routes and predictions from the model with starting material information. The compounds that are missing from the prediction are in red.
-Our model correctly reproduces both routes for Mitapivat (Fig.
+This underscores that the performance of our models relies upon sufficient representation of transformations in the training set.
+This limitation, however, should naturally disappear once larger multistep datasets become available.
+Inability of DMS-Flash to suggest reduction of amide carbonyl in Fig.
+5 was the motivation for the creation of mixture-of-experts (MoE) models.
+We conjectured that different experts might specialize in different classes of transformations, and even if one expert dominates the generations, one could still manually force the output from a different expert.
+Fortunately, no such manual interventions were needed as DMS-Flex (Duo) predicted the route almost identical to the second transformation as the Top-1 candidate (Fig.
+6).
+The only difference from the correct route is the order of transformations: in the correct route the tosyl group is introduced after amidation with methylamine, not before.
+Figure 6: A variation of second Vonoprazan route predicted by DMS-Flex (Duo).
+The correct route is shown in Fig.
+5 and differs by the order of the last steps: tosyl group is introduced after amidation, not before.
+Figure 7: Daridorexant literature multistep routes and predictions from DMS-Flex (Duo). The compounds that are missing from the prediction are in dotted boxes.
+DMS-Flex (Duo) correctly reproduces both routes for Mitapivat (Fig.
 5 and Fig.
@@
 In contrast, the route for Daridorexant (Fig.
-6) is only partially predicted: our model does not split the right child of the root node (Daridorexant itself) into the two compounds highlighted in red.
+7) is only partially predicted: our model does not split the right child of the root node into the two compounds in the dotted boxes.
 All other intermediates and starting materials are predicted correctly.
-The reason for this behavior is not well understood, but it can be rationalized by looking at the distribution of the number of leaves at root nodes: 73% of root nodes (target compounds) have at least 1 leaf (see SI Fig.
-7).
-Such dependence on the distributions of the training set comprises an important limitation of the multi-step first approach.
-This issue would not occur with methods based on single-step prediction because the tree search framework would not halt at compounds outside the stock compound set before reaching the maximum iterations.
-Ensuring that multi-step first models reproduce that aspect of SSR methods constitutes a direction for future research.
+The reason for this behavior is not well understood, but it can be rationalized by looking at the distribution of the number of leaves at root nodes: 73% of root nodes (target compounds) have at least 1 leaf (see Fig.
+S3).
+This dependence on the distributions of the training set comprises an important limitation of our models; however, it should also be resolved once bigger datasets become available.
+Admittedly, this issue would not occur with current methods based on single-step prediction because the tree search framework would not halt at compounds outside the stock compound set before reaching the maximum iterations.
+Ensuring that multistep-first models reproduce that aspect of SSR methods constitutes a direction for future research.
+This example, however, does not imply that the model struggles to predict routes with convergent branches.
+If one defines a convergent route as the one that has at least one node (compound) with at least two children that are not leaves, and repeats the analysis, no significant differences in Top-K accuracies (Fig.
+S5-S6) or in the number of routes found (Fig.
+S7-S8) is observed despite significant variation in the number of routes that are convergent among training routes, n1, and n5 (Fig.
+S3).
 4 Conclusion
-In this work, we introduced DirectMultiStep, a transformer-based model that directly generates multi-step synthetic routes for a given target compound.
+In this work, we introduce the first-of-its-kind inherently multistep approach for multistep retrosynthesis prediction.
 By predicting the entire route as a single string, our approach bypasses the need for iterative single-step predictions and exponential search space traversal.
-The model accommodates specific conditions such as the desired number of steps and starting materials, enabling efficient route planning.
-DirectMultiStep outperforms state-of-the-art methods on the PaRoutes dataset, achieving a 2.2x improvement in Top-1 accuracy on the n1 test set and a 3.3x improvement on the n5 test set.
-The model also successfully predicts routes for FDA-approved drugs not included in the training data, showcasing its generalization capabilities.
-However, the model’s performance may be impacted by the suboptimal diversity of the current training set, particularly for less common reaction types.
-Despite these limitations, the DirectMultiStep approach presents a promising direction towards fully automated retrosynthetic planning.
-As larger and more diverse multi-step datasets become available, the model’s performance is expected to improve further.
-Future research should focus on addressing the limitations identified, such as the need to split longer routes into sub-routes and the dependence on the distributions of the training set.
+We present a family of transformer-based DirectMultiStep (DMS) models which can predict correct routes just from the structure of the target compound, but can also accommodate specific conditions such as the desired number of steps and starting materials, enabling efficient route planning.
+Our flagman DMS-Flex (Duo) model, incorporating mixture-of-experts approach, outperforms state-of-the-art methods on the PaRoutes dataset, achieving a 2.5x improvement in Top-1 accuracy on the n1 test set and a 3.9x improvement on the n5 test set.
+We demonstrate the quality of predicted routes by testing our models on several FDA-approved drugs.
+Those predictions also highlight limitations caused by the suboptimal diversity of the current training set, particularly for less common reaction types.
+Those limitations are partially mitigated by employing mixture-of-experts techniques, making them a promising candidate for further research in retrosynthesis prediction and applications of generative models to chemistry in general.
+We believe our work demonstrates the superiority of multistep-first approach to full route prediction and inspires further research into methods that reduce the exponential complexity of iterative application of single-step based methods.
+We are grateful to the authors of PaRoutes for the development and release of the first multistep prediction benchmark, and we hope more benchmarks based on experimentally verified routes become available.
+Future research may focus on improvement of performance on longer routes (e.g.
+eliminating the necessity to split longer routes) and (or) incorporation of additional constraints (such as cost) on the route prediction.
 Overall, this work demonstrates the potential of transformer-based models for efficient and scalable computer-aided synthesis planning.
-5 Broad Impact
-Chemical synthesis of target compounds of interest forms the routine for millions of organic chemists around the world, including graduate students in PhD programs and synthetic chemists in pharmaceutical companies.
-The ability to find shorter (and thus more efficient) routes to any desired compound is an enticing challenge with the potential of saving on labor costs and chemical waste.
-As a result, multi-step prediction methods have the potential to improve the lives of millions of people, and potentially contribute to faster advances in the development of new drugs and materials.
-We hope that our work constitutes an important advancement in that direction.
-Quite unfortunately, however, there is no fundamental chemical difference between how safe and useful compounds (such as drugs) and malign compounds (such as chemical weapons) are synthesized.
-Both require application of the same set of rules of organic synthesis, and so the ability to make drugs faster and more efficiently inherently brings along the ability to make poisons faster and more efficiently.
-One cannot impose restrictions on the tokens (chemical elements) learned by the model, as even some combinations of nitrogen and oxygen (which are widespread in proteins and DNA) may be cytotoxic.
-Therefore, unless the humanity at large decides that the benefits of automated retrosynthesis do not outweigh its potential harms, the most we can do is communicate the dangers clearly and perhaps impose restrictions that would still require a trained (which often accompanies training in ethics) chemist to use our model.
-Besides, all meta-data of reaction conditions and catalysts are removed from the predictions of our model, which would present a deliberate barrier against malicious uses.
-6 Code and Data Availability
+We release the code used for training and route prediction as a public GitHub Repo under MIT License.
+Our models are also available with a Web GUI at models.batistalab.com.
+5 More Computational Details
+5.1 Data Curation and Preparation
+A second version of the PaRoutes [44] dataset containing 450k routes is stripped of all metadata and stored as recursive dictionaries representing multistep routes.
+The two evaluation sets, n1 and n5, are processed similarly.
+A training partition is created by removing all repetitions (some routes were represented multiple times with different metadata) and permutations (swapping left and right subtrees) of routes in n1 and n5 from the full dataset, resulting in 163k routes.
+The training dataset is augmented by adding 2 permutations for each route.
+To train the DMS model which takes starting material as an input, we find all starting materials (leaves) for each tree and store a combination of the target compound with each starting material as a separate entry.
+As a result, DMS with SM is trained on 1 340 243 inputs, and the DMS without SM on 432 684 inputs.
+All SMILES strings are tokenized by treating each character as a token, and the string representation of the multistep route is tokenized similarly, treating delimiters of the tree (‘smiles’, ‘children’, ,, [, ], {, and }) as separate tokens.
+The final vocabulary size is 52 (including start, end, and padding tokens), the largest multistep route has 1074 tokens, while the largest target compound and starting material have 145 and 135 tokens, respectively.
+5.2 Model Architecture
+We present a family of DirectMultiStep (DMS) models with varying parameter counts and architectural details.
+DMS-Flash and DMS-Deep the classical transformer encoder-decoder architecture [56], whereas DMS-Flex and DMS-Wide incorporate sparsely-gated mixture-of-experts layers [46, 47, 48].
+Major parameters choices are given in model card (Table  1).
+All models use GeLU activations [57].
+Intermediate experiments with evaluations on subsets of n1 and n5 with 50 and 500 routes reveal no benefit in increasing the number of hidden dimensions to 512 (Tables S4-S5), which could be rationalized by noticing that 256 is already almost 5x the vocabulary size.
+Although a feed-forward multiplier of four is more common in NLP models, we find no significant improvement in performance as compared to the feed-forward multiplier of three (Tables S6-S7).
+In addition, the number of encoder layers does not have to be the same as the number of decoder layers, and models with 6 encoder and 12 decoder layers (trained with SM) performed almost as good as the models with 12 encoder and decoder layers.
+For MoE models, we find no significant improvement upon increase of number of experts to four, and models trained with two active experts outperform those trained with one.
+5.3 Training
+During training, a 10% dropout rate is used in embedding, attention, and feed-forward layers.
+The AdamW optimizer [58, 2019adamw] is used with a learning rate scheduler that warms up from 0 to $3.5\times 10^{-4}$ for DMS-Flex and DMS-Wide, $3\times 10^{-4}$ for DMS-Flash, and $1\times 10^{-4}$ for DMS-Deep during the first 10% of training steps and undergoes cosine decay to $10\%$ of the maximal learning rate over 20 epochs.
+The optimal initial learning rate for each model was chosen by starting from 5e-4 and monitoring for gradient explosion (in which case the learning rate was decreased by 0.5e-4).
+Gradients are clipped to 1.0.
+In addition, each token in the input to the encoders has a 5% chance of being masked.
+During hyperparameter optimization, the model was validated on a 5% split from the training routes (which exclude n1 and n5 routes).
+6 Data and Software Availability
 The reaction route dataset is from the 2.0 version of PaRoutes in their GitHub repository [44], available under the Apache License 2.0.
 Code to process dataset, implementation of model architecture, code for training, generation, and evaluation are available under MIT License at https://github.com/batistagroup/DirectMultiStep.
-7 Acknowledgments
+A web interface is also available at models.batistalab.com.
+Author Contributions
+YS, HL, and AM contributed equally to this work.
+YS conceived the idea.
+YS and HL conducted preliminary model training and testing.
+AM introduced the mixture of experts (MoE) approach and carried out additional model training and testing.
+VSB provided guidance for the research.
+All authors participated in writing and reviewing the manuscript.
+Conflict of Interest
+The authors declare no conflict of interest.
+7 Acknowledgement
 The authors acknowledge a generous allocation of high-performance computing time from NERSC. The development of the methodology was supported by the NSF CCI grant (VSB, Award Number 2124511).
+Supporting Information
+Figure S1: Lengths of valid routes predicted by DMS-Explorer for targets in n1 and n5 compared to the length of the correct (experimental) route.
+Left: distribution of differences in lengths for all predictions for all targets.
+Right: distribution of average (among all predictions for a single target) differences.
+Figure S2: Lengths of valid routes predicted by DMS-Flex (Duo) for targets in n1 and n5 compared to the length of the correct (experimental) route.
+Left: distribution of differences in lengths for all predictions for all targets.
+Right: distribution of average (among all predictions for a single target) differences.
+Figure S3: Distribution of the relative frequencies for the number of leaves (nodes with no children) at root node.
+Table S1: Number of predicted routes in n1 that pass post-processing filters.
+Model Valida Correct SMb,f SMs In Stockc Correct and In Stockd,f
+Flash 9998 9942 8814 8725
+Flashe 9994 7660
+Explorer 9997 9923 8681 8555
+Explorere 9995 7418
+Explorer-XLe 9998 8008
+Flex (Mono) 9941 9545 8338 7948
+Flex (Duo) 9995 9866 8828 8649
+Flex (Duo)e 9994 7692
+Deepe 9998 7957
+Widee 9996 8089
+aAt least one route with all valid SMILES (including target, intermediates, and SMs).
+bAt least one route with the correct conditioned SMs (and with all valid SMILES).
+cAt least one route with all SMs in the n1 stock set (and with all valid SMILES).
+dAt least one route with the correct conditioned SMs and with all SMs in the n1 stock set (and with all valid SMILES).
+eSMs not provided during inference.
+fBlank indicates not relevant (models without SM conditions).
+Table S2: Number of predicted routes in n5 that pass post-processing filters.
+Model Valida Correct SMb,f SMs In Stockc Correct and In Stockd,f
+Flash 9991 9888 8646 8487
+Flashe 9991 7576
+Explorer 9997 9903 8439 8281
+Explorere 9998 7227
+Explorer-XLe 10000 7904
+Flex (Mono) 9932 9374 8037 7528
+Flex (Duo) 9993 9808 8627 8382
+Flex (Duo)e 9990 7605
+Deepe 9997 7830
+Widee 9997 7950
+aAt least one route with all valid SMILES (including target, intermediates, and SMs).
+bAt least one route with the correct conditioned SMs (and with all valid SMILES).
+cAt least one route with all SMs in the n5 stock set (and with all valid SMILES).
+dAt least one route with the correct conditioned SMs and with all SMs in the n5 stock set (and with all valid SMILES).
+eSMs not provided during inference.
+fBlank indicates not relevant (models without SM conditions).
+Table S3: Specifications of architecture for different model variants.
+DMS Model Encoder Decoder Hidden FF Total Active Experts
+Variant Layers Layers Dimension Multipliera Experts Training Generation
+Van-19 6 6 256 3 1 1 1
+Van-60 8 8 512 4 1 1 1
+MoE-1/3 6 6 256 3 3 1 1
+a an integer by which the hidden dimension is multiplied in feed-forward layers after attention.
+Figure S4: Fraction of routes (grouped by route length) that are convergent in training set, n1, and n5.
+A route is considered convergent if any node has at least two children that are not leaves.
+Overall, convergent routes constitute 9.4% of the training routes, 5.6% of n1, and 10.9% of n5.
+Figure S5: Top-K accuracy on the n1 test set for different models for (a) all routes in n1, (b) convergent routes in n1, and (c) non-convergent routes in n1.
+A route is considered convergent if any node has at least two children that are not leaves.
+Figure S6: Top-K accuracy on the n5 test set for different models for (a) all routes in n5, (b) convergent routes in n5, and (c) non-convergent routes in n5.
+A route is considered convergent if any node has at least two children that are not leaves.
+Figure S7: Distribution of the average number of routes for targets in n1 test set (a-c) after checking for the validity of smiles, (d-f) after canonicalization of SMILES and removing repetitions (considering permutations of children), and (g-i) after filtering for routes that have all reactants in stock.
+Distribution is shown separately for all targets (a, d, g), for targets for which an experimental route is convergent (b, e, h) and for targets for which an experimenal route is non-convergent (c, f, i).
+A route is considered convergent if any node has at least two children that are not leaves.
+Figure S8: Distribution of the average number of routes for targets in n5 test set (a-c) after checking for the validity of smiles, (d-f) after canonicalization of SMILES and removing repetitions (considering permutations of children), and (g-i) after filtering for routes that have all reactants in stock.
+Distribution is shown separately for all targets (a, d, g), for targets for which an experimental route is convergent (b, e, h) and for targets for which an experimenal route is non-convergent (c, f, i).
+A route is considered convergent if any node has at least two children that are not leaves.
+Table S4: Top-$K$ accuracy on route test set-n1 (10 000 routes).
+DMS SM SM
+Models Size Steps (Train) (Gen.) Top-1 Top-2 Top-3 Top-4 Top-5 Top-10
+Van-19 19M 0.44 0.53 0.57 0.59 0.60 0.63
+MoE-1/3 38M 0.40 0.47 0.50 0.51 0.52 0.53
+Van-60 60M 0.38 0.46 0.48 0.50 0.51 0.52
+Van-60 60M 0.36 0.42 0.44 0.45 0.45 0.46
+Table S5: Top-$K$ accuracy on route test set-n5 (10 000 routes).
+DMS SM SM
+Models Size Steps (Train) (Gen.) Top-1 Top-2 Top-3 Top-4 Top-5 Top-10
+Van-19 19M 0.39 0.47 0.50 0.52 0.53 0.56
+MoE-1/3 38M 0.36 0.42 0.45 0.46 0.46 0.48
+Van-60 60M 0.38 0.46 0.48 0.50 0.51 0.52
+Van-60 60M 0.36 0.42 0.44 0.45 0.45 0.46
+Table S6: Top-$K$ accuracy on subseta of 500 routes from test set-n1.
+SM SM
+Models Size Steps (Train) (Gen.) Top-1 Top-2 Top-3 Top-4 Top-5 Top-10
+Flash 10M 0.402 0.476 0.522 0.550 0.564 0.582
+6x4-12x4b 18M 0.446 0.552 0.598 0.616 0.624 0.640
+12x3-12x3 19M 0.452 0.570 0.596 0.618 0.626 0.636
+12x4-12x4 22M 0.452 0.552 0.588 0.610 0.620 0.642
+6x3-6x3 10M 0.292 0.362 0.384 0.388 0.396 0.402
+aThe subset was chosen randomly with a seed of 42. Flash results are provided to establish how the performance on these 500 routes relates to the performance on full test set.
+bThe model names follow NxF-KxF format, where N is the number of encoder layers, K is the number of decoder layers, and F is the feed-forward multiplier
+Table S7: Top-$K$ accuracy on subseta of 500 routes from test set-n5.
+SM SM
+Models Size Steps (Train) (Gen.) Top-1 Top-2 Top-3 Top-4 Top-5 Top-10
+Flash 10M 0.374 0.442 0.474 0.484 0.498 0.516
+6x4-12x4b 18M 0.400 0.464 0.498 0.520 0.530 0.548
+12x3-12x3 19M 0.406 0.472 0.496 0.518 0.524 0.538
+12x4-12x4 22M 0.370 0.448 0.472 0.498 0.508 0.526
+6x3-6x3 10M 0.222 0.264 0.282 0.288 0.290 0.298
+aThe subset was chosen randomly with a seed of 42. Flash results are provided to establish how the performance on these 500 routes relates to the performance on full test set.
+bThe model names follow NxF-KxF format, where N is the number of encoder layers, K is the number of decoder layers, and F is the feed-forward multiplier
+Table S8: Top-1 Accuracy on n1 routes split by route length
+Model 2 3 4 5 6 7 8 9 10
+Flash 0.538 0.410 0.323 0.273 0.329 0.308 0.286 0.400 0.000
+Explorera 0.480 0.346 0.282 0.266 0.305 0.323 0.286 0.500 0.000
+Explorer-XLa,b 0.402 0.304 0.271 0.259 0.296 0.277 0.286 0.600 0.000
+Flex (Mono) 0.456 0.318 0.226 0.193 0.183 0.154 0.143 0.300 0.000
+Flex (Duo) 0.554 0.426 0.348 0.292 0.362 0.354 0.286 0.500 0.000
+Flex (Duo)b 0.471 0.360 0.310 0.278 0.319 0.277 0.286 0.400 0.000
+Deepb 0.497 0.395 0.338 0.268 0.272 0.262 0.286 0.600 0.000
+Wideb 0.511 0.416 0.344 0.308 0.329 0.292 0.286 0.600 0.000
+aSteps not provided during inference.
+bSMs not provided during inference.
+Table S9: Top-10 Accuracy on n1 routes split by route length
+Top10 accuracy on routes with length
+Model 2 3 4 5 6 7 8 9 10
+Flash 0.763 0.600 0.465 0.385 0.408 0.400 0.333 0.500 0.000
+Explorera 0.692 0.546 0.417 0.373 0.376 0.431 0.333 0.500 0.000
+Explorer-XLa,b 0.608 0.463 0.387 0.363 0.390 0.400 0.333 0.600 0.000
+Flex (Mono) 0.650 0.474 0.325 0.254 0.286 0.231 0.143 0.300 0.000
+Flex (Duo) 0.765 0.610 0.469 0.373 0.408 0.385 0.333 0.600 0.000
+Flex (Duo)b 0.594 0.460 0.388 0.336 0.371 0.338 0.333 0.600 0.000
+Deepb 0.641 0.519 0.442 0.364 0.362 0.323 0.333 0.600 0.000
+Wideb 0.647 0.531 0.434 0.378 0.390 0.354 0.333 0.600 0.000
+aSteps not provided during inference.
+bSMs not provided during inference.
+Table S10: Top-1 Accuracy on n5 routes split by route length
+Top1 accuracy on routes with length
+Model 2 3 4 5 6 7 8 9 10
+Flash 0.502 0.401 0.337 0.320 0.331 0.328 0.429 0.692 0.333
+Explorera 0.458 0.350 0.300 0.289 0.343 0.288 0.286 0.462 0.333
+Explorer-XLa,b 0.395 0.317 0.284 0.303 0.313 0.248 0.405 0.692 0.333
+Flex (Mono) 0.436 0.324 0.258 0.211 0.191 0.184 0.190 0.154 0.333
+Flex (Duo) 0.549 0.416 0.366 0.336 0.352 0.328 0.476 0.615 0.333
+Flex (Duo)b 0.465 0.358 0.324 0.312 0.326 0.344 0.429 0.615 0.333
+Deepb 0.474 0.384 0.343 0.313 0.338 0.272 0.381 0.385 0.333
+Wideb 0.477 0.399 0.352 0.337 0.366 0.304 0.381 0.615 0.333
+aSteps not provided during inference.
+bSMs not provided during inference.
+Table S11: Top-10 Accuracy on n5 routes split by route length
+Top10 accuracy on routes with length
+Model 2 3 4 5 6 7 8 9 10
+Flash 0.712 0.583 0.479 0.431 0.425 0.400 0.429 0.692 0.333
+Explorera 0.654 0.529 0.442 0.404 0.434 0.376 0.405 0.692 0.333
+Explorer-XLa,b 0.586 0.469 0.416 0.404 0.423 0.384 0.47 0.692 0.333
+Flex (Mono) 0.621 0.461 0.357 0.285 0.297 0.272 0.333 0.462 0.333
+Flex (Duo) 0.721 0.594 0.489 0.436 0.444 0.416 0.500 0.692 0.333
+Flex (Duo)b 0.581 0.462 0.407 0.383 0.421 0.400 0.452 0.692 0.333
+Deepb 0.604 0.507 0.451 0.414 0.430 0.368 0.429 0.615 0.333
+Wideb 0.616 0.524 0.451 0.420 0.432 0.328 0.476 0.692 0.333
+aSteps not provided during inference.
+bSMs not provided during inference.

### v3

@@
 Traditional computer-aided synthesis planning (CASP) methods rely on iterative single-step predictions, leading to exponential search space growth that limits efficiency and scalability.
-We introduce a series of transformer-based models, utilizing mixture of experts approach, that directly generate multistep synthetic routes as a single string by conditionally predicting each molecule based on all preceding ones.
-Our models can accommodate specific conditions such as the desired number of steps and starting materials, with the top-performing DMS-Flex (Duo) surpassing state-of-the-art methods on the PaRoutes dataset with a 2.5x improvement in Top-1 accuracy on the n1 test set and a 3.9x improvement on the n5 test set.
-It also successfully predicts routes for FDA-approved drugs not included in the training data, showcasing its generalization capabilities.
-While the current suboptimal diversity of the training set may impact performance on less common reaction types, our multistep-first approach presents a promising direction towards fully automated retrosynthetic planning.
+We introduce a series of transformer-based models, that leverage a mixture of experts approach to directly generate multistep synthetic routes as a single string, conditionally predicting each transformation based on all preceding ones.
+Our DMS Explorer XL model, which requires only target compounds as input, outperforms state-of-the-art methods on the PaRoutes dataset with 1.9x and 3.1x improvements in Top-1 accuracy on the n1 and n5 test sets, respectively.
+Providing additional information, such as the desired number of steps and starting materials, enables both a reduction in model size and an increase in accuracy, highlighting the benefits of incorporating more constraints into the prediction process.
+The top-performing DMS-Flex (Duo) model scores 25-50% higher on Top-1 and Top-10 accuracies for both n1 and n5 sets.
+Additionally, our models successfully predict routes for FDA-approved drugs not included in the training data, demonstrating strong generalization capabilities.
+While the limited diversity of the training set may affect performance on less common reaction types, our multistep-first approach presents a promising direction towards fully automated retrosynthetic planning.
 1 Introduction
@@
 The number of steps is the largest number of reactions from SM to the target compound (tree height).
+[…]
+[…]
+Figure 1: The workflow of DirectMultiStep.
+(a) The SMILES representation of the target compound (blue), starting material (red, optional), and the number of steps (optionally) are tokenized, concatenated, and fed into our transformer model.
+The model predicts a string representation of the multistep synthesis tree.
+Spaces are added for clarity, and indentations indicate the levels in the synthesis route (tree).
+(b) Molecular structures corresponding to the target compound (blue), starting material (red, optional), and the predicted synthesis tree with structures of all molecules.
 2.2 Route Representation
@@
 Our models predict that string by taking SMILES of the target compound, (optionally) starting material, and (optionally) number of steps as input.
-[…]
-[…]
-Figure 1: The workflow of DirectMultiStep.
-(a) The SMILES representation of the target compound (blue), (optional) starting material (red), and the number of steps are tokenized, concatenated, and fed into our transformer model.
-The model predicts a string representation of the multistep synthesis tree.
-Spaces are added for clarity, and indentations indicate the levels in the synthesis route (tree).
-(b) Molecular structures corresponding to the target compound (blue), (optional) starting material (red), and the predicted synthesis tree with structures of all molecules.
-2.3 Evaluation Metrics and Post-Processing
-Evaluation is performed on n1 and n5 (10 000 routes each) evaluation sets from PaRoutes [44], which have been created by selecting the most diverse routes from the USPTO.
-Fig.
-2 shows the distribution of route lengths in training set, n1, and n5.
-The distribution in both n1 and n5 is shifted towards longer routes; for example, the fraction of routes with 4 steps or more is 15% for training set, 27% for n1 and 51% for n5.
 Figure 2: Distribution of the relative frequencies of route lengths (in terms of number of steps) in the training dataset before augmentation with permutations (163 689 routes, black), n1 test set (10 000 routes, blue), and n5 test set (10 000 routes, purple).
 Distribution is split into routes shorter (left subplot) and longer than 6 steps (right subplot).
-The structure of target compound, starting material (optionally, the deepest leaf of the correct route is chosen during evaluation), and number of steps is passed to the transformer encoder.
+2.3 Route Generation and Post-Processing
+The structure of target compound, starting material (optionally), and number of steps (optionally) is passed to the transformer encoder.
 A beam search with a width of 50 is employed to generate 50 candidate routes by sampling next tokens from the decoder.
@@
 This adjustment ensures that shorter routes with simpler compounds are not disproportionately favored while maintaining a balance between route length and overall sequence probability.
-After beam search finds end-of-sequence tokens for all 50 candidates, predicted routes are checked for validity of all SMILES strings (which are canonicalized), the presence of starting materials in the stock set (following procedure of PaRoutes taken to be the set of all leaves in n1 or n5, and absence of repetitions (which can be present as permutations of other routes).
+After beam search finds end-of-sequence tokens for all 50 candidates, predicted routes are checked for validity of all SMILES strings (which are canonicalized), the presence of starting materials in the stock set, and absence of repetitions (which can be present as permutations of other routes).
 Remaining routes are compared against the true (experimental) route.
@@
 presence of correct route ranked $K$ or lower.
-Importantly, the Top-$K$ accuracy is evaluated against all 10 000 targets, not just the ones for which routes are found.
-2.4 Models Presented
-As a baseline, we train a conventional transformer-based sequence-to-sequence model with 6 encoder and 6 decoder layers (DMS-Flash) with structures of starting materials provided during training.
-To achieve comparable performance without starting material as a hint, a significantly larger model with 12 encoder and 36 decoder layers is needed (DMS-Deep).
-We also present the first, to our knowledge, application of mixture-of-experts (MoE [46, 47, 48]) architecture to a chemical problem.
-To create MoE model, one replaces the 2-layer MLP that follows every attention block with an ensemble of 2-layer MLPs (called experts).
-A preceding router determines (for each input) the $n$-most relevant (active) experts, and passes the input through each of them.
-Output of each active expert is added with weights determined by the router.
-Notably, the number of experts active during inference does not have to be equal to the number of experts active during training, although the latter tends to be optimal.
-We present two MoE models, the DMS-Flex (trained with SM) and DMS-Wide (trained without SM).
-The relevant architectural details of models are given in Table 1.
-Table 1: Specifications of architecture for different model variants.
-DMS Model Encoder Decoder Hidden FF Total Active Experts
-Variant Layers Layers Dimension Multipliera Experts Training Generation
-Flash 6 6 256 3 1 1 1
-Explorer 6 6 256 3 3 2 2
-Explorer XL 6 24 256 3 3 2 2
-Flex (Mono) 6 6 256 3 3 2 1
-Flex (Duo) 6 6 256 3 3 2 2
-Deep 12 36 256 3 1 1 1
-Wide 12 12 256 3 3 2 2
-a an integer by which the hidden dimension is multiplied in feed-forward layers after attention.
 3 Results and Discussion
-3.1 Top-K Accuracy on Evaluation Sets
-3.1.1 Vanilla Transformers
-Top-$K$ accuracy is shown in Table 2 for set-n1 and Table 3 for set-n5.
-The smallest DMS-Flash model (which receives structure of SM during generation) achieves state-of-the-art performance, surpassing all previous methods in all Top-$K$ accuracies, with the most significant increase in Top-1 accuracy: 2.4x on n1 and 3.7x on n5.
-To achieve comparable performance without providing the structure of the SM, a significantly larger model DMS-Deep is needed.
-This establishes a trade-off between size of the model (and thus generation speed) and the amount of information it receives.
-The limitations of the necessity to provide the structure of the SM are discussed in Section 3.1.3.
-Quite remarkably, a Flash model trained with structures of SM may generate routes even when SMs are not provided (Tables 2-3).
-While that results in a 20% drop in Top-1 accuracy, it is significantly better than a model with the same architecture that is trained without SM (Tables S6-S7).
-In other words, providing structure of a correct starting material helps the model learn better and improves performance even if the starting material is not used during evaluation.
+3.1 Overview
+In this work, we present DMS Explorer XL, a mixture-of-experts based model (50M parameters), that predicts full multistep retrosynthetic route given the structure of the target compound.
+Our model shows comparable search performance on ChemBL-5000[46] and significantly improves Top-1 accuracy on n1 and n5 evaluation datasets from PaRoutes[44].
+We find that if one is willing to provide extra information to the encoder, such as desired route length and (or) starting material, the model size can be reduced and simultaneously both Top-1 and Top-10 accuracy can be increased significantly.
+We present a family of such models (model cards are in Sec.
+5.2) and discuss their use cases.
+3.2 Search Performance
+First, we report the solved rate on ChemBL-5000 [46] (Table 1).
+Our baseline DMS Explorer XL (50M) shows comparable performance to that of a tree search on predictions from a single-step model (AiZynthFinder [46]).
+A slightly smaller DMS Wide (40M) model that was trained with desired route length as an extra parameter to the encoder shows significantly better performance.
+Despite 7x more calls to the model (with different route length parameters), the average run time per target is only 3x higher.
+Notably, a 4x smaller DMS Flash (10M) model trained with both route length and structure of the starting material, can be evaluated on new targets even without the structure of the starting material with a minor (3-5%) decrease in the solved rate, but at the same cost as the DMS Explorer XL model.
+Table 1: Search performance for DirectMultiStep models on ChemBL-5000 [46]
+Methoda Stock Setb Solved Rate Run Time (s)c
+DMS Explorer XL Buyables 68.66% 26.9
+AiZynthFinder [46] In-House + Enamine 75.24% NA
+DMS Explorer XL eMolecules Screening 75.58% 26.9
+DMS Flashd,e Buyables 86.66% 23.1
+DMS Flashd,e eMolecules Screening 88.54% 23.1
+DMS Wided Buyables 91.20% 75.4
+DMS Wided eMolecules Screening 91.94% 75.4
+a All DMS models are run with a beam size of 50 on a single NVIDIA A100 GPU with half-precision floating point inference (FP16).
+b eMolecules Screening is from Chen et al.
+[39] and contains 23.1M screening compounds from eMolecules as of 2019 (2019-11-01).
+Buyables is from Roh et al.
+[47] and includes 0.329M buyable building blocks from eMolecules, Sigma-Aldrich, Mcule, ChemBridge Hit2Lead, and WuXi LabNetwork.
+In-House + Enamine is from Genheden et al.
+[46] and is not publicly available.
+c Averages over all targets. NA indicates that the run time is not available in the original publication.
+d Uses step counts from 2 to 8 (total of 7 DMS model runs).
+e SMs not provided for DMS-Flash.
+Next, we report the solved rate on n1 and n5 evaluation sets (each contains 10 000 routes) in Table 2.
+Following the procedure of PaRoutes [44] the starting material stock set is taken to be the set of all leaves in n1 or n5.
+DMS Explorer XL performs noticeably worse than MCTS or Retro*, but scores better than DFPN on n5.
+The number and the distribution of solved targets before enforcing the stock set is shown in Tables S1-S2 and Figures S1-S2.
+Other DMS models that accept desired route length (Wide) and also the structure of the starting material (Flash) solve more targets, but still underperform Retro*.
+Unfortunately, the training set of PaRoutes contains routes that include 47 of the targets from the commonly used USPTO-190 evaluation subset, which makes direct comparison with existing models unfair.
+For completeness, we report solved rate on all 190 and 143 (excluding the 47) targets in Table S3.
+Table 2: Search performance for DirectMultiStep models on n1 and n5 evaluation sets
+Methoda Evaluation Set Solved Targets First Solutionb Search Time (s)b
+DMS-Explorer-XL n1 8008 NA 14.7
+DMS-Wide n1 8089 NA 6.0
+DFPN [44] n1 8475 43.0 347.3
+DMS-Flash n1 8814 NA 2.0
+MCTS [44] n1 9714 8.6 303.3
+Retro* [44] n1 9726 7.0 300.7
+DFPN [44] n5 7382 53.2 297.9
+DMS-Explorer-XL n5 7904 NA 16.3
+DMS-Wide n5 7950 NA 7.4
+DMS-Flash n5 8646 NA 2.5
+MCTS [44] n5 9676 11.7 365.7
+Retro* [44] n5 9703 10.5 349.2
+a All DMS models are run with a beam size of 50 on a single NVIDIA A100 GPU with half-precision floating point inference (FP16).
+b Averages over all targets.
+3.3 Top-K Accuracy (Route Quality)
+An attentive reader might notice the absence of a measure of a feasibility of proposed transformations during the discussion of solved rate above.
+One might argue that the number of solved targets only matters if the proposed transformations to the set of commercially available materials can be performed experimentally.
+In the absence of a cheap ab initio tool to predict experimental feasibility of the proposed transformation, following Genheden et al.
+[44] and Liu et al.
+[43], we utilize Top-K accuracy on reproducibility of experimentally verified routes as the proxy for route quality.
 Table 3: Top-$K$ accuracy on route test set-n1 (10 000 routes).
@@
 Models Size Steps (Train) (Gen.) Top-1 Top-2 Top-3 Top-4 Top-5 Top-10
+Explorer-XL 50M 0.32 0.40 0.44 0.45 0.46 0.48
+Deep 41M 0.36 0.44 0.47 0.49 0.50 0.52
+Explorer 19M 0.36 0.45 0.48 0.51 0.52 0.55
+Explorer 19M 0.27 0.33 0.36 0.37 0.39 0.40
 Flash 10M 0.41 0.50 0.54 0.56 0.58 0.60
 Flash 10M 0.32 0.39 0.42 0.44 0.45 0.47
-Explorer 19M 0.36 0.45 0.48 0.51 0.52 0.55
-Explorer 19M 0.27 0.33 0.36 0.37 0.39 0.40
-Explorer-XL 50M 0.32 0.40 0.44 0.45 0.46 0.48
+Wide 38M 0.38 0.46 0.49 0.50 0.51 0.53
 Flex (Mono) 19M 0.32 0.40 0.43 0.44 0.45 0.48
@@
 Flex (Duo) 19M 0.34 0.41 0.44 0.45 0.46 0.47
-Deep 41M 0.36 0.44 0.47 0.49 0.50 0.52
-Wide 38M 0.38 0.46 0.49 0.50 0.51 0.53
 aThese data are collected from the 2.0 version of PaRoutes in their GitHub repository [44].
@@
 Models Size Steps (Train) (Gen.) Top-1 Top-2 Top-3 Top-4 Top-5 Top-10
+Explorer-XL 50M 0.31 0.38 0.41 0.42 0.43 0.45
+Deep 41M 0.33 0.40 0.43 0.45 0.46 0.48
+Explorer 19M 0.33 0.40 0.43 0.45 0.46 0.49
+Explorer 19M 0.25 0.31 0.34 0.35 0.36 0.37
 Flash 10M 0.37 0.44 0.47 0.49 0.51 0.53
 Flash 10M 0.30 0.36 0.39 0.40 0.41 0.43
-Explorer 19M 0.33 0.40 0.43 0.45 0.46 0.49
-Explorer 19M 0.25 0.31 0.34 0.35 0.36 0.37
-Explorer-XL 50M 0.31 0.38 0.41 0.42 0.43 0.45
+Wide 38M 0.35 0.42 0.44 0.46 0.47 0.49
 Flex (Mono) 19M 0.28 0.34 0.36 0.38 0.39 0.41
@@
 Flex (Duo) 19M 0.32 0.38 0.40 0.42 0.42 0.44
-Deep 41M 0.33 0.40 0.43 0.45 0.46 0.48
-Wide 38M 0.35 0.42 0.44 0.46 0.47 0.49
 aThese data are collected from the 2.0 version of PaRoutes in their GitHub repository [44].
-Remarkably, the Top-2 accuracy of DMS-Flash and the Top-4 accuracy of DMS-Deep exceed the Top-10 accuracy of all search methods that utilize iterative application of single-step models, highlighting the advantages of the multistep-first approach.
-To understand this performance difference, it helps to notice that iterative application of single-step calls results in a breadth-first search-like exploration of a node, as all reactants leading to a compound are generated at the same time.
+We report Top-$K$ accuracy on n1 and n5 sets, each containing 10 000 routes.
+Notably, both sets were constructed to maximize internal diversity (see the PaRoutes paper [44] for the procedure), and as seen in Fig.
+2, differ significantly from the training set in route length distribution.
+The DMS Explorer XL model (only target required) shows significant improvement on Top-1 accuracy for both n1 (1.8x, Table 3) and n5 (3.1x, Table 4) sets, while matching (n1) or improving (n5) performance on Top-10 accuracy.
+We attribute this improvement to the multistep-first nature of our approach.
+To understand the distinction from existing approaches, it helps to notice that iterative application of single-step calls results in a breadth-first search-like exploration of a node: all reactants leading to a compound are generated at the same time.
 Our approach, however, is a depth-first search-like: the whole left sub-tree of target compound in Fig.
@@
 This provides an extra context for the prediction of the second precursor to the target compound (or any other right child), which reduces the search space and finds more relevant compounds faster.
-A few words on the relevance of Top-$K$ accuracy as the primary evaluation metric are in order.
-One might question whether the number of targets for which valid (i.e., having valid SMILES) routes are found (or solvability rate if all SMs are in a given stock set) and the average length of routes predicted should be considered.
-It is essential to realize that even if you ensure that every single-step transformation corresponds to a known reaction by template matching, the route may still be experimentally not feasible simply because atoms that are unchanged (i.e., the ones ignored by the template) contribute to the energies of the reacting orbitals through inductive and resonance effects, thereby changing the energetic favorability of the reaction.
-As a result, short of performing highly-accurate ab initio calculations of energies (after corresponding geometry optimization) with all relevant solvation effects and considering thermal corrections, there is no computational metric that would assess experimental feasibility of the predicted route.
-In addition, one could argue that the ability to discover an experimentally verified route is a prerequisite for finding shorter, but still feasible, routes.
-Therefore, Top-$K$ accuracy is the closest proxy to route quality.
-Additionally, the ability to predict correct routes with better ranks is a significant improvement because it reduces the number of candidates to be passed through the aforementioned highly expensive energy calculations.
-Finally, one could argue that single-step methods do not provide inherent ranking of candidates, so the performance of search methods may depend on the choice of ranking metric.
-However, that is yet another benefit of our approach: there is no need in devising another algorithm or model for route ranking, the beam probabilities are sufficiently accurate.
 The superior performance of DMS models does not incur additional computational cost.
-A search with a beam size of 50 takes 3-16 seconds (depending on the number of steps) with a DMS-Flash or a DMS-Flex model on a single GPU (NVIDIA M4000), which is comparable to the time required to find the first successful route using previous methods (7-50 seconds [44]).
+A search with a beam size of 50 takes 1-40 seconds (depending on the complexity of the target and model size) on a single NVIDIA A100 GPU, which is comparable to the time required to find the first successful route using previous methods (7-50 seconds) [44].
+3.3.1 Input information and model size trade-off
+A unifying pattern of results in Tables 3-4 is that if one is willing to provide more information to the model, model size can be reduced and performance can be improved at the same time.
+For example, providing the desired route length reduces model size by 20% and improves accuracy by 2-4% (DMS Deep).
+The distribution of generated route lengths compared to the length of the experimental route is shown in Fig.
+S3-S4.
+Alternatively, providing the structure of the starting material (SM) allows to reduce the parameter count in half, and the resulting DMS Explorer (19M) model scores 4-7% better on all Top-$K$ accuracies.
+Providing desired route length can further half parameter count (DMS Flash is only 10M) and add another 4-5% in Top-1 and Top-10 accuracy.
+Notably, models that take extra information during training might still be used without that extra information, and would perform better than models of the same size trained without extra information.
+For example, a DMS Flash is trained with SM but might be evaluated without the structure of the SM, and the resulting performance (Tables 3-4) is 4-7% better than that of equivalent 6-layer transformer trained without SM (entry 6x3-6x3 in Tables S4-S5).
+These results suggest a new training strategy if one is constrained by computational resources: extra input, derived directly from training data (DMS Flash is trained with the structure of the deepest node provided as SM) might serve as a learning aid, even if not used in route generation in production.
+3.3.2 Mixture of experts architecture reduces cost of generation
+Route generation time depends not only on the total number of parameters, but also on how those parameters are distributed.
+DMS Deep (38 M) utilizes a vanilla 36-layer decoder, meaning that every token prediction has to sequentially pass through 36 attention and multi-layer perceptron blocks.
+We find that redistributing the model parameters from the number of layers into the width of each layer by employing mixture-of-experts [48, 49, 50] based architecture (see Sec.
+5.2 for details) allows a threefold reduction in the number of layers (and thus a threefold speedup): DMS Wide (38M) needs only 12-layer decoder and performs 1-2% better than DMS Deep.
+Another important advantage of MoE-based models is that tokens passed through different active experts can be processed in parallel, which speeds up route generation even further.
+Preliminary experiments showed that decoders with fewer than 6 layers perform significantly worse, making it impossible to train a size-comparable MoE alternative to DMS Flash.
+However, a 6-layer MoE trained with structures of SM (DMS Flex, 19M) similarly outperforms DMS Flash by 1-2% at no increase in average runtime given that extra parameters are processed in parallel expert blocks.
+Just like DMS Flash, DMS Flex can be evaluated without SM and still have better Top-1 accuracy on both n1/n5 and better Top-10 on n5 than existing methods.
+Interestingly, MoE-based models provide another dimension for flexibility: routes can be generated with a number of active experts different than that used during training.
+While DMS Flex (trained with 2 active experts) evaluated with 2 active experts (Duo mode) shows higher Top-K accuracy than if evaluated with 1 active expert (Mono mode), both result in comparable number of routes at different stages of post-processing (Tables S1-S2).
+As a result, one could consider predictions from Mono mode as a source of diversity.
+Our experimentation with the increase in the number of experts, the number of encoder/decoder layers, or the feed-forward multiplier evaluated on small subsets of n1 or n5, show no significant improvement in performance (Tables S4-S8), which most likely means that the 20 M (with SM) or 40 M (without SM) are optimal given the current size of the training set.
 Figure 3: Distribution of Top-1 and Top-10 accuracy of predictions with DMS-Flex (Duo) on test sets n1 and n5.
@@
 4
-3.1.2 MoE Transformers
-The DMS-Flex (Flexible), a mixture-of-experts model trained with two out of three experts active, provides the largest performance improvement when evaluated with two experts active (Duo mode): 2.5x increase in Top-1 on n1 and 3.9x on n5.
-Evaluating DMS-Flex with one expert active (Mono) mode worsens the performance by 30%; given that the number of valid routes (Tables S1-S2) does not change significantly, one could consider the Mono mode as a source of diversity in predictions.
-Notably, even though the DMS-Flex is trained with SM, it can be evaluated without providing SM (empty padding tokens are fed into the encoder instead) which still results in 34% Top-1 accuracy and 47+Top-10.
-Evaluating DMS-Flex without providing the number of steps as an input drops performance significantly, but a different model with the same architecture trained without number of steps as an input, DMS-Explorer (Fig.
-S1 shows the distribution of route lengths predicted by Explorer relative to the length of an experimental route), results in only a 8% drop in accuracy.
-Using a larger DMS-Wide model trained without SM provides comparable performance to that of DMS-Deep.
-An important advantage of DMS-Wide is that tokens passed through different active experts can be processed on a GPU in parallel, which enables faster predictions.
-In addition, DMS-Wide has fewer layers, so fewer consecutive calculations have to be performed; in practice, doubling the number of encoder layers results in doubling of generation time.
-One might also conclude that similar performance of DMS-Deep and DMS-Wide, which have comparable parameter counts, implies that the expressive power mostly depends on the number of parameters, and not on the extra non-linearities provided by larger number of layers.
-Our experimentation with the increase in the number of experts, the number of encoder/decoder layers, or the feed-forward multiplier evaluated on small subsets of n1 or n5, show no significant improvement in performance (Tables S4-S7), which most likely means that the 20 M (with SM) or 40 M (without SM) are optimal given the current size of the training set.
+3.3.3 Route quality as a function of route length
 Fig.
@@
 However, the performance on routes with 5-8 steps is comparable to that of shorter routes, and the performance on 9-step routes is even comparable to that of 2-step routes.
-The exceptional accuracy on 9-step routes could be rationalized by the fact that their relative abundance in n1 and n5 does not exceed the relative abundance in the training partition (unlike for routes with length 3-7, as seen in Fig.
-3).
-Top-1 and Top-10 accuracies of other DMS models for different route lengths are given in Tables S8-S11.
+Top-1 and Top-10 accuracies of other DMS models for different route lengths are given in Tables S9-S12.
 […]
@@
 (b) Correct prediction from DMS-Flex (Duo) for the second half of the 10-step route with starting material information.
-3.1.3 Use Cases for Different DMS Model Variants
-The requirement to provide the starting material structure as input to the model might, in theory, seem as a limitation, limiting the relevance of DMS-Flash/Flex models.
-However, in practice, a chemist often can know the structure of a desired starting material because Corey’s retrosynthetic framework allows finding at least one route to starting materials by performing one functional group transformation at a time.
-Such deterministically found routes will often be too long to be efficient in terms of yield (overall yield is the product of yields of individual steps).
-The art of organic synthesis lies in finding ways to perform several transformations simultaneously using a minimal number of protecting groups.
-As such, we believe our SM-based models provide value as an assistant to this task: finding better ways to get to the target compound from a starting material.
-The requirement to specify the number of steps as an input becomes a useful option in this context.
-Moreover, if the target compound is chiral (i.e., it has a mirror image which cannot be superimposed), even during manual retrosynthetic planning, chemists check if chiral centers can be obtained from chiral starting materials (such as naturally occurring amino acids) because developing reactions that create chiral centers in the desired conformation remains an active research area (and subject of 2001 and 2021 Nobel Prizes in Chemistry).
-An example of such a route is shown in Fig.
-7.
-In addition, one can use more computationally expensive DMS-Wide/Deep models to discover potential starting materials, and then use those structures as inputs to DMS-Flash/Flex to generate suggestions for each starting material of interest.
-To lift the necessity to specify desired route length, a user also has a choice between running Flex/Wide models several times with different desired route lengths or simply running Explorer-XL once.
-Of course, ideally one would prefer a single model that predicts the whole route with just from the structure of the target compound, without number of steps or starting material provided as an extra input.
-We believe better performance could be attainable in the future once bigger datasets become available.
-Finding ways to improve performance even with current dataset constitutes an important area of further research.
-3.2 Separation of Long Routes into Shorter Routes
+3.3.4 Separation of Long Routes into Shorter Routes
 As seen in Fig.
@@
 4).
+3.3.5 Use Cases for Different DMS Model Variants
+The requirement to provide the starting material structure as an input to the model might, in theory, seem as a limitation, limiting the relevance of DMS Flash/Flex models.
+However, in practice, a chemist often can know the structure of a desired starting material because Corey’s retrosynthetic framework allows finding at least one route to starting materials by performing one functional group transformation at a time.
+Such deterministically found routes will often be too long to be efficient in terms of yield (overall yield is the product of yields of individual steps).
+The art of organic synthesis lies in finding ways to perform several transformations simultaneously using a minimal number of protecting groups.
+As such, we believe our SM-based models provide value as an assistant to this task: finding better ways to get to the target compound from a starting material.
+The requirement to specify the number of steps as an input becomes a useful option in this context.
+Moreover, if the target compound is chiral (i.e., it has a mirror image which cannot be superimposed), even during manual retrosynthetic planning, chemists check if chiral centers can be obtained from chiral starting materials (such as naturally occurring amino acids) because developing reactions that create chiral centers in the desired conformation remains an active research area (and subject of 2001 and 2021 Nobel Prizes in Chemistry).
+An example of such a route is shown in Fig.
+7.
+Alternatively, one can use more computationally expensive DMS Wide/Deep models to discover potential starting materials, and then use those structures as inputs to DMS Flash/Flex to generate suggestions for each starting material of interest.
+To lift the necessity to specify desired route length, a user also has a choice between running Flex/Wide models several times with different desired route lengths or simply running Explorer XL once.
+Of course, ideally one would prefer a single model that predicts the whole route with just from the structure of the target compound, without number of steps or starting material provided as an extra input.
+We believe better performance could be attainable in the future once bigger datasets become available.
+Finding ways to improve performance even with current dataset constitutes an important area of further research.
 […]
@@
 The reason for this behavior is not well understood, but it can be rationalized by looking at the distribution of the number of leaves at root nodes: 73% of root nodes (target compounds) have at least 1 leaf (see Fig.
-S3).
+S5).
 This dependence on the distributions of the training set comprises an important limitation of our models; however, it should also be resolved once bigger datasets become available.
@@
 This example, however, does not imply that the model struggles to predict routes with convergent branches.
-If one defines a convergent route as the one that has at least one node (compound) with at least two children that are not leaves, and repeats the analysis, no significant differences in Top-K accuracies (Fig.
-S5-S6) or in the number of routes found (Fig.
+If one defines a convergent route as the one that has at least one node (compound) with at least two children that are not leaves, and repeats the analysis, no significant differences in the number of routes found (Fig.
+S1-S2) or Top-K accuracies (Fig.
 S7-S8) is observed despite significant variation in the number of routes that are convergent among training routes, n1, and n5 (Fig.
-S3).
+S6).
 4 Conclusion
@@
 We present a family of transformer-based DirectMultiStep (DMS) models which can predict correct routes just from the structure of the target compound, but can also accommodate specific conditions such as the desired number of steps and starting materials, enabling efficient route planning.
-Our flagman DMS-Flex (Duo) model, incorporating mixture-of-experts approach, outperforms state-of-the-art methods on the PaRoutes dataset, achieving a 2.5x improvement in Top-1 accuracy on the n1 test set and a 3.9x improvement on the n5 test set.
+Our flagman DMS-Flex (Duo) model, incorporating mixture-of-experts approach, outperforms state-of-the-art methods on the PaRoutes dataset, achieving a 2.5x improvement in Top-1 accuracy on an n1 test set and a 3.9x improvement on the n5 test set.
 We demonstrate the quality of predicted routes by testing our models on several FDA-approved drugs.
@@
 We present a family of DirectMultiStep (DMS) models with varying parameter counts and architectural details.
-DMS-Flash and DMS-Deep the classical transformer encoder-decoder architecture [56], whereas DMS-Flex and DMS-Wide incorporate sparsely-gated mixture-of-experts layers [46, 47, 48].
-Major parameters choices are given in model card (Table  1).
+DMS-Flash and DMS-Deep employ the classical transformer encoder-decoder architecture [58], whereas DMS-Flex and DMS-Wide incorporate sparsely-gated mixture-of-experts layers [48, 49, 50].
+To create MoE model, one replaces the 2-layer MLP that follows every attention block with an ensemble of 2-layer MLPs (called experts).
+A preceding router determines (for each input) the $n$-most relevant (active) experts, and passes the input through each of them.
+Output of each active expert is added with weights determined by the router.
+Notably, the number of experts active during inference does not have to be equal to the number of experts active during training, although the latter tends to be optimal.
+Major parameters choices are given in model card (Table  5).
 All models use GeLU activations [59].
-Intermediate experiments with evaluations on subsets of n1 and n5 with 50 and 500 routes reveal no benefit in increasing the number of hidden dimensions to 512 (Tables S4-S5), which could be rationalized by noticing that 256 is already almost 5x the vocabulary size.
-Although a feed-forward multiplier of four is more common in NLP models, we find no significant improvement in performance as compared to the feed-forward multiplier of three (Tables S6-S7).
+Table 5: Specifications of architecture for different model variants.
+DMS Model Encoder Decoder Hidden FF Total Active Experts
+Variant Layers Layers Dimension Multipliera Experts Training Generation
+Flash 6 6 256 3 1 1 1
+Explorer 6 6 256 3 3 2 2
+Explorer XL 8 24 256 3 3 2 2
+Flex (Mono) 6 6 256 3 3 2 1
+Flex (Duo) 6 6 256 3 3 2 2
+Deep 12 36 256 3 1 1 1
+Wide 12 12 256 3 3 2 2
+a an integer by which the hidden dimension is multiplied in feed-forward layers after attention.
+Intermediate experiments with evaluations on subsets of n1 and n5 with 50 and 500 routes reveal no benefit in increasing the number of hidden dimensions to 512 (Tables S4-S6), which could be rationalized by noticing that 256 is already almost 5x the vocabulary size.
+Although a feed-forward multiplier of four is more common in NLP models, we find no significant improvement in performance as compared to the feed-forward multiplier of three (Tables S7-S8).
 In addition, the number of encoder layers does not have to be the same as the number of decoder layers, and models with 6 encoder and 12 decoder layers (trained with SM) performed almost as good as the models with 12 encoder and decoder layers.
@@
 During training, a 10% dropout rate is used in embedding, attention, and feed-forward layers.
-The AdamW optimizer [58, 2019adamw] is used with a learning rate scheduler that warms up from 0 to $3.5\times 10^{-4}$ for DMS-Flex and DMS-Wide, $3\times 10^{-4}$ for DMS-Flash, and $1\times 10^{-4}$ for DMS-Deep during the first 10% of training steps and undergoes cosine decay to $10\%$ of the maximal learning rate over 20 epochs.
+The AdamW optimizer [60, 61] is used with a learning rate scheduler that warms up from 0 to $3.5\times 10^{-4}$ for DMS-Flex and DMS-Wide, $3\times 10^{-4}$ for DMS-Flash, and $1\times 10^{-4}$ for DMS-Deep during the first 10% of training steps and undergoes cosine decay to $10\%$ of the maximal learning rate over 20 epochs.
 The optimal initial learning rate for each model was chosen by starting from 5e-4 and monitoring for gradient explosion (in which case the learning rate was decreased by 0.5e-4).
@@
 Author Contributions
-YS, HL, and AM contributed equally to this work.
+YS and AM contributed equally to this work.
 YS conceived the idea.
 YS and HL conducted preliminary model training and testing.
-AM introduced the mixture of experts (MoE) approach and carried out additional model training and testing.
+AM introduced the mixture of experts (MoE) approach.
+AM and YS carried out additional model training and testing and wrote the manuscript.
 VSB provided guidance for the research.
-All authors participated in writing and reviewing the manuscript.
+All authors have given approval to the final version of the paper.
 Conflict of Interest
@@
 Supporting Information
-Figure S1: Lengths of valid routes predicted by DMS-Explorer for targets in n1 and n5 compared to the length of the correct (experimental) route.
-Left: distribution of differences in lengths for all predictions for all targets.
-Right: distribution of average (among all predictions for a single target) differences.
-Figure S2: Lengths of valid routes predicted by DMS-Flex (Duo) for targets in n1 and n5 compared to the length of the correct (experimental) route.
-Left: distribution of differences in lengths for all predictions for all targets.
-Right: distribution of average (among all predictions for a single target) differences.
-Figure S3: Distribution of the relative frequencies for the number of leaves (nodes with no children) at root node.
 Table S1: Number of predicted routes in n1 that pass post-processing filters.
 Model Valida Correct SMb,f SMs In Stockc Correct and In Stockd,f
+Explorer-XLe 9998 8008
+Deepe 9998 7957
+Explorer 9997 9923 8681 8555
+Explorere 9995 7418
 Flash 9998 9942 8814 8725
 Flashe 9994 7660
-Explorer 9997 9923 8681 8555
-Explorere 9995 7418
-Explorer-XLe 9998 8008
+Widee 9996 8089
 Flex (Mono) 9941 9545 8338 7948
@@
 Flex (Duo)e 9994 7692
-Deepe 9998 7957
-Widee 9996 8089
 aAt least one route with all valid SMILES (including target, intermediates, and SMs).
@@
 Model Valida Correct SMb,f SMs In Stockc Correct and In Stockd,f
+Explorer-XLe 10000 7904
+Deepe 9997 7830
+Explorer 9997 9903 8439 8281
+Explorere 9998 7227
 Flash 9991 9888 8646 8487
 Flashe 9991 7576
-Explorer 9997 9903 8439 8281
-Explorere 9998 7227
-Explorer-XLe 10000 7904
+Widee 9997 7950
 Flex (Mono) 9932 9374 8037 7528
@@
 Flex (Duo)e 9990 7605
-Deepe 9997 7830
-Widee 9997 7950
 aAt least one route with all valid SMILES (including target, intermediates, and SMs).
@@
 fBlank indicates not relevant (models without SM conditions).
-Table S3: Specifications of architecture for different model variants.
-DMS Model Encoder Decoder Hidden FF Total Active Experts
-Variant Layers Layers Dimension Multipliera Experts Training Generation
-Van-19 6 6 256 3 1 1 1
-Van-60 8 8 512 4 1 1 1
-MoE-1/3 6 6 256 3 3 1 1
-a an integer by which the hidden dimension is multiplied in feed-forward layers after attention.
-Figure S4: Fraction of routes (grouped by route length) that are convergent in training set, n1, and n5.
-A route is considered convergent if any node has at least two children that are not leaves.
-Overall, convergent routes constitute 9.4% of the training routes, 5.6% of n1, and 10.9% of n5.
-Figure S5: Top-K accuracy on the n1 test set for different models for (a) all routes in n1, (b) convergent routes in n1, and (c) non-convergent routes in n1.
-A route is considered convergent if any node has at least two children that are not leaves.
-Figure S6: Top-K accuracy on the n5 test set for different models for (a) all routes in n5, (b) convergent routes in n5, and (c) non-convergent routes in n5.
-A route is considered convergent if any node has at least two children that are not leaves.
-Figure S7: Distribution of the average number of routes for targets in n1 test set (a-c) after checking for the validity of smiles, (d-f) after canonicalization of SMILES and removing repetitions (considering permutations of children), and (g-i) after filtering for routes that have all reactants in stock.
+Figure S1: Distribution of the average number of routes predicted by DMS Explorer XL for targets in n1 test set (a-c) after checking for the validity of smiles, (d-f) after canonicalization of SMILES and removing repetitions (considering permutations of children), and (g-i) after filtering for routes that have all reactants in stock.
 Distribution is shown separately for all targets (a, d, g), for targets for which an experimental route is convergent (b, e, h) and for targets for which an experimenal route is non-convergent (c, f, i).
 A route is considered convergent if any node has at least two children that are not leaves.
-Figure S8: Distribution of the average number of routes for targets in n5 test set (a-c) after checking for the validity of smiles, (d-f) after canonicalization of SMILES and removing repetitions (considering permutations of children), and (g-i) after filtering for routes that have all reactants in stock.
+Figure S2: Distribution of the average number of routes predicted by DMS Explorer XL for targets in n5 test set (a-c) after checking for the validity of smiles, (d-f) after canonicalization of SMILES and removing repetitions (considering permutations of children), and (g-i) after filtering for routes that have all reactants in stock.
 Distribution is shown separately for all targets (a, d, g), for targets for which an experimental route is convergent (b, e, h) and for targets for which an experimenal route is non-convergent (c, f, i).
 A route is considered convergent if any node has at least two children that are not leaves.
-Table S4: Top-$K$ accuracy on route test set-n1 (10 000 routes).
-DMS SM SM
-Models Size Steps (Train) (Gen.) Top-1 Top-2 Top-3 Top-4 Top-5 Top-10
-Van-19 19M 0.44 0.53 0.57 0.59 0.60 0.63
-MoE-1/3 38M 0.40 0.47 0.50 0.51 0.52 0.53
-Van-60 60M 0.38 0.46 0.48 0.50 0.51 0.52
-Van-60 60M 0.36 0.42 0.44 0.45 0.45 0.46
-Table S5: Top-$K$ accuracy on route test set-n5 (10 000 routes).
-DMS SM SM
-Models Size Steps (Train) (Gen.) Top-1 Top-2 Top-3 Top-4 Top-5 Top-10
-Van-19 19M 0.39 0.47 0.50 0.52 0.53 0.56
-MoE-1/3 38M 0.36 0.42 0.45 0.46 0.46 0.48
-Van-60 60M 0.38 0.46 0.48 0.50 0.51 0.52
-Van-60 60M 0.36 0.42 0.44 0.45 0.45 0.46
+Table S3: Search performance for DirectMultiStep models on USPTO-190 [39]
+Methoda Stock Setb Solved Ratec Run Time (s)d
+DMS-Explorer-XL eMolecules Screening 33.2% (31.5%) 21.9
+MCTS [39] eMolecules Screening 33.7% 370.5
+DMS-Flashe,f eMolecules Screening 60.5% (55.2%) 31.3
+DMS-Widee eMolecules Screening 63.7% (56.6%) 105.8
+Retro* [39] eMolecules Screening 86.8% 156.6
+RetroGraph [41] eMolecules Screening 99.5% NA
+DMS-Explorer-XL Buyables 27.9% (28.0%) 21.9
+Original + MCTS [47] Buyables 46.3% NA
+DMS-Flashe,f Buyables 55.3% (51.0%) 31.3
+DMS-Widee Buyables 56.8% (51.7%) 105.8
+Higherlev + MCTS [47] Buyables 73.7% NA
+a All DMS models are run with a beam size of 50 on a single NVIDIA A100 GPU with half-precision floating point inference (fp16).
+b eMolecules Screening is from Chen et al.
+[39] and contains 23.1M screening compounds from eMolecules as of 2019 (2019-11-01).
+Buyables is from Roh et al.
+[47] and includes 0.329M buyable building blocks from eMolecules, Sigma-Aldrich, Mcule, ChemBridge Hit2Lead, and WuXi LabNetwork.
+c 47 of the targets from USPTO-190 are present as non-leaf nodes in PaRoutes training set. The solved rate on the remaining 143 targets is reported in parentheses.
+d Averages over all targets. NA indicates that the run time is not available in the original publication.
+e Uses step counts from 2 to 8 (total of 7 DMS model runs).
+f SMs not provided for DMS-Flash.
+Figure S3: Lengths of valid routes predicted by DMS-Explorer for targets in n1 and n5 compared to the length of the correct (experimental) route.
+Left: distribution of differences in lengths for all predictions for all targets.
+Right: distribution of average (among all predictions for a single target) differences.
+Figure S4: Lengths of valid routes predicted by DMS-Flex (Duo) for targets in n1 and n5 compared to the length of the correct (experimental) route.
+Left: distribution of differences in lengths for all predictions for all targets.
+Right: distribution of average (among all predictions for a single target) differences.
 Table S4: Top-$K$ accuracy on subseta of 500 routes from test set-n1.
@@
 bThe model names follow NxF-KxF format, where N is the number of encoder layers, K is the number of decoder layers, and F is the feed-forward multiplier
+Table S6: Specifications of architecture for different model variants.
+DMS Model Encoder Decoder Hidden FF Total Active Experts
+Variant Layers Layers Dimension Multipliera Experts Training Generation
+Van-19 6 6 256 3 1 1 1
+Van-60 8 8 512 4 1 1 1
+MoE-1/3 6 6 256 3 3 1 1
+a an integer by which the hidden dimension is multiplied in feed-forward layers after attention.
+Table S7: Top-$K$ accuracy on route test set-n1 (10 000 routes).
+DMS SM SM
+Models Size Steps (Train) (Gen.) Top-1 Top-2 Top-3 Top-4 Top-5 Top-10
+Van-19 19M 0.44 0.53 0.57 0.59 0.60 0.63
+MoE-1/3 38M 0.40 0.47 0.50 0.51 0.52 0.53
+Van-60 60M 0.38 0.46 0.48 0.50 0.51 0.52
+Van-60 60M 0.36 0.42 0.44 0.45 0.45 0.46
+Table S8: Top-$K$ accuracy on route test set-n5 (10 000 routes).
+DMS SM SM
+Models Size Steps (Train) (Gen.) Top-1 Top-2 Top-3 Top-4 Top-5 Top-10
+Van-19 19M 0.39 0.47 0.50 0.52 0.53 0.56
+MoE-1/3 38M 0.36 0.42 0.45 0.46 0.46 0.48
+Van-60 60M 0.38 0.46 0.48 0.50 0.51 0.52
+Van-60 60M 0.36 0.42 0.44 0.45 0.45 0.46
 Table S9: Top-1 Accuracy on n1 routes split by route length
 Model 2 3 4 5 6 7 8 9 10
+Explorer-XLa,b 0.402 0.304 0.271 0.259 0.296 0.277 0.286 0.600 0.000
+Deepb 0.497 0.395 0.338 0.268 0.272 0.262 0.286 0.600 0.000
+Explorera 0.480 0.346 0.282 0.266 0.305 0.323 0.286 0.500 0.000
 Flash 0.538 0.410 0.323 0.273 0.329 0.308 0.286 0.400 0.000
-Explorera 0.480 0.346 0.282 0.266 0.305 0.323 0.286 0.500 0.000
-Explorer-XLa,b 0.402 0.304 0.271 0.259 0.296 0.277 0.286 0.600 0.000
+Wideb 0.511 0.416 0.344 0.308 0.329 0.292 0.286 0.600 0.000
 Flex (Mono) 0.456 0.318 0.226 0.193 0.183 0.154 0.143 0.300 0.000
@@
 Flex (Duo)b 0.471 0.360 0.310 0.278 0.319 0.277 0.286 0.400 0.000
-Deepb 0.497 0.395 0.338 0.268 0.272 0.262 0.286 0.600 0.000
-Wideb 0.511 0.416 0.344 0.308 0.329 0.292 0.286 0.600 0.000
 aSteps not provided during inference.
@@
 Model 2 3 4 5 6 7 8 9 10
+Explorer-XLa,b 0.608 0.463 0.387 0.363 0.390 0.400 0.333 0.600 0.000
+Deepb 0.641 0.519 0.442 0.364 0.362 0.323 0.333 0.600 0.000
+Explorera 0.692 0.546 0.417 0.373 0.376 0.431 0.333 0.500 0.000
 Flash 0.763 0.600 0.465 0.385 0.408 0.400 0.333 0.500 0.000
-Explorera 0.692 0.546 0.417 0.373 0.376 0.431 0.333 0.500 0.000
-Explorer-XLa,b 0.608 0.463 0.387 0.363 0.390 0.400 0.333 0.600 0.000
+Wideb 0.647 0.531 0.434 0.378 0.390 0.354 0.333 0.600 0.000
 Flex (Mono) 0.650 0.474 0.325 0.254 0.286 0.231 0.143 0.300 0.000
@@
 Flex (Duo)b 0.594 0.460 0.388 0.336 0.371 0.338 0.333 0.600 0.000
-Deepb 0.641 0.519 0.442 0.364 0.362 0.323 0.333 0.600 0.000
-Wideb 0.647 0.531 0.434 0.378 0.390 0.354 0.333 0.600 0.000
 aSteps not provided during inference.
@@
 Model 2 3 4 5 6 7 8 9 10
+Explorer-XLa,b 0.395 0.317 0.284 0.303 0.313 0.248 0.405 0.692 0.333
+Deepb 0.474 0.384 0.343 0.313 0.338 0.272 0.381 0.385 0.333
+Explorera 0.458 0.350 0.300 0.289 0.343 0.288 0.286 0.462 0.333
 Flash 0.502 0.401 0.337 0.320 0.331 0.328 0.429 0.692 0.333
-Explorera 0.458 0.350 0.300 0.289 0.343 0.288 0.286 0.462 0.333
-Explorer-XLa,b 0.395 0.317 0.284 0.303 0.313 0.248 0.405 0.692 0.333
+Wideb 0.477 0.399 0.352 0.337 0.366 0.304 0.381 0.615 0.333
 Flex (Mono) 0.436 0.324 0.258 0.211 0.191 0.184 0.190 0.154 0.333
@@
 Flex (Duo)b 0.465 0.358 0.324 0.312 0.326 0.344 0.429 0.615 0.333
-Deepb 0.474 0.384 0.343 0.313 0.338 0.272 0.381 0.385 0.333
-Wideb 0.477 0.399 0.352 0.337 0.366 0.304 0.381 0.615 0.333
 aSteps not provided during inference.
@@
 Model 2 3 4 5 6 7 8 9 10
+Explorer-XLa,b 0.586 0.469 0.416 0.404 0.423 0.384 0.47 0.692 0.333
+Deepb 0.604 0.507 0.451 0.414 0.430 0.368 0.429 0.615 0.333
+Explorera 0.654 0.529 0.442 0.404 0.434 0.376 0.405 0.692 0.333
 Flash 0.712 0.583 0.479 0.431 0.425 0.400 0.429 0.692 0.333
-Explorera 0.654 0.529 0.442 0.404 0.434 0.376 0.405 0.692 0.333
-Explorer-XLa,b 0.586 0.469 0.416 0.404 0.423 0.384 0.47 0.692 0.333
+Wideb 0.616 0.524 0.451 0.420 0.432 0.328 0.476 0.692 0.333
 Flex (Mono) 0.621 0.461 0.357 0.285 0.297 0.272 0.333 0.462 0.333
@@
 Flex (Duo)b 0.581 0.462 0.407 0.383 0.421 0.400 0.452 0.692 0.333
-Deepb 0.604 0.507 0.451 0.414 0.430 0.368 0.429 0.615 0.333
-Wideb 0.616 0.524 0.451 0.420 0.432 0.328 0.476 0.692 0.333
 aSteps not provided during inference.
 bSMs not provided during inference.
+Figure S5: Distribution of the relative frequencies for the number of leaves (nodes with no children) at root node.
+Figure S6: Fraction of routes (grouped by route length) that are convergent in training set, n1, and n5.
+A route is considered convergent if any node has at least two children that are not leaves.
+Overall, convergent routes constitute 9.4% of the training routes, 5.6% of n1, and 10.9% of n5.
+Figure S7: Top-K accuracy on the n1 test set for different models for (a) all routes in n1, (b) convergent routes in n1, and (c) non-convergent routes in n1.
+A route is considered convergent if any node has at least two children that are not leaves.
+Figure S8: Top-K accuracy on the n5 test set for different models for (a) all routes in n5, (b) convergent routes in n5, and (c) non-convergent routes in n5.
+A route is considered convergent if any node has at least two children that are not leaves.